### v1

# Super-resolving frequency measurement with mode-selective quantum memory

Abstract
High-precision optical frequency measurement is indispensable to modern science and technology, yet conventional spectroscopic techniques struggle to resolve spectral features narrower than their bandwidth. Here, we introduce a unique platform for super-resolving frequency estimation utilizing a mode-selective atomic Raman quantum memory implemented in warm cesium vapor. By precisely engineering the light-matter interaction, our memory coherently stores the optimal temporal mode with high fidelity and retrieves it on-demand, realizing a mode crosstalk as low as 0.34%. For the task of estimating the separation between two spectral lines, we experimentally measure the mean squared error of the frequency estimate, achieving a sensitivity as small as 1/20 of the linewidth with up to a (34±$\pm 4$)-fold enhancement in precision over direct intensity measurements. This enhanced resolution of frequency measurements, combined with the memory’s capability of on-demand storage, retrieval and mode conversion, paves the way for multifunctional memory-based time-frequency sensors and their integration within quantum networks.
I Introduction
The time-frequency (TF) degree of freedom of light underpins a broad range of applications spanning high-resolution spectroscopy [1], precision timekeeping [2, 3], ultrafast optics [4] and emerging quantum technologies [5, 6, 7, 8]. In spectroscopy, for example, resolving spectral lines is essential for probing the atomic and molecular properties of matter. Frequency measurements are also crucial for quantum metrology and sensing [9, 10, 11], as they largely rely on measuring the transition between energy levels of quantized physical systems. However, the precision of these optical measurements has long been constrained by instrumental limitations. Every spectral feature has an intrinsic linewidth set by the Fourier limit - a lower limit being the inverse of either the measurement time or the signal’s temporal duration. The resolving power of conventional spectrometers is compounded by the Rayleigh criterion [12], which defines the minimal resolvable separation between two spectral lines. As the line separation approaches zero, the uncertainty in their resolution escalates rapidly. This phenomenon, often termed Rayleigh’s curse, poses a critical barrier, particularly when probing weak or photon‐limited signals.
The ultimate precision in resolving spectral features is bounded by the quantum mechanical nature of the optical field [13]. From quantum estimation theory, Tsang, Nair, and Lu have shown that Rayleigh’s criterion is not a fundamental limit, but rather an artifact of the direct intensity measurement (DI) strategy - with coherent measurements an arbitrarily small separation can be resolved with a constant, fine precision [14]. To circumvent the classical resolution limit, implementing a coherent mode filter to select the optimal mode basis before detection, instead of DI measurements, is essential [15, 16, 17, 18, 19, 20]. The practical realization of this advantage, however, relies critically on the accuracy of the mode filtering, as experimental imperfections like mode crosstalk and detector noise substantially diminish the achievable precision [21, 22, 23, 24, 25, 26]. The high-fidelity coherent mode filtering is generally needed for broader quantum information processing and quantum metrology technologies [6, 7, 27, 28, 29]. Looking ahead, the future quantum networks demand sensor nodes with such capability of coherent TF processing, ideally integrated with on-demand buffering for signal synchronization [30, 31], coherent bandwidth and frequency conversion to interface different physical platforms and network channels [32, 33, 34] - all while preserving quantum coherence. These integrated functionalities are essential for building robust, distributed quantum-enhanced sensors [35] capable of dynamically adjusting to varying environmental conditions across a network [36, 37, 38, 39].
In TF super-resolution, quantum pulse gates (QPGs), which use nonlinear waveguides for mode-selective frequency conversion, have been deployed to resolve temporal and spectral separations for ultrafast pulses with hundreds of GHz bandwidth [40, 41]. While QPGs demonstrate programmable mode selectivity [42], they inherently lack on-demand storage and buffering capabilities. In the narrow-band frequency regime, time-inversion interferometry using gradient echo memory has demonstrated sub-Rayleigh resolution for tens of kHz bandwidth pulses [43], but they are limited in the ability to select arbitrary modes and require cumbersome magnetic field gradients and cryogenics. Additionally, there have also been works focusing on frequency super-resolution across tens of GHz bandwidth using dispersion engineering and electro-optic modulators [44, 45]. However, no existing platform, including conventional spectroscopic tools, delivers high‐precision super‐resolution in the MHz to GHz bandwidth, while also providing on‐demand storage, retrieval and user-defined mode selectivity.
In this work, we introduce a new scheme of TF super-resolution via high-fidelity coherent mode filtering in atomic Raman quantum memory. Photonic quantum memories have been widely studied as devices capable of absorbing and re-emitting photonic states on demand, with applications in quantum networks, communication and computing [46, 47, 48, 49, 50, 51, 52]. Implemented in a warm cesium vapor, our Raman memory utilizes a three-level atomic system where a strong control field coherently maps an incoming weak signal field onto a collective atomic spin-wave coherence via stimulated Raman absorption. By tailoring the temporal profile of the Raman control pulses, we demonstrate high-fidelity mode-selectivity up to 99.6% for orthogonal Hermite-Gaussian (HG) temporal modes. With this high-fidelity, user-defined coherent mode filter, our memory-based platform stores the optimal signal mode containing the information of spectral line separation and performs detection on the retrieved mode. We calibrate the measurement apparatus and extract frequency separations via a maximum-likelihood estimation procedure applied to the retrieved mode’s photon statistics. Focusing on sub-linewidth frequency separation under various detected photon budgets (from $2\times 10^{3}$ to $100\times 10^{3}$), our platform consistently outperforms DI methods, achieving significant precision enhancements. Operating in the MHz to GHz bandwidth, our memory-based platform extends the toolbox of TF metrology and provides integrated functionalities encompassing mode filtering, buffering, and shaping. These capabilities are ideally suited for the development of next-generation quantum sensor nodes and their integration into future quantum networks.
II Results
II.1 Super-resolving measurement in frequency domain
The task of resolving two closely spaced spectral features can be framed as the estimation of a single parameter: the normalized frequency separation $\epsilon=\Delta\omega/\sigma$, where $\Delta\omega$ denotes the frequency difference between the two spectral lines centered at $\omega_{0}\pm\Delta\omega/2$, and $\sigma$ is the spectral linewidth. We consider the signal to originate from two mutually incoherent emitters of equal intensity, each characterized by a Gaussian spectral amplitude $\psi(\omega)=(2\pi\sigma^{2})^{-1/4}\exp(-\omega^{2}/4\sigma^{2})$. Conventional spectroscopy measures the power spectrum, characterized by a normalized intensity profile given by
$S(\omega|\epsilon)=\frac{1}{2}\left[\left|\psi\left(\omega-\omega_{0}-\frac{\epsilon\sigma}{2}\right)\right|^{2}+\left|\psi\left(\omega-\omega_{0}+\frac{\epsilon\sigma}{2}\right)\right|^{2}\right].$
The separation $\epsilon$ is then estimated by fitting this model to the measured spectrum. However, as $\epsilon\to 0$, the two Gaussian lines increasingly overlap, making it exceedingly difficult to distinguish them and leading to a large estimation error. For this DI strategy, the amount of information about $\epsilon$ extractable from the measurement outcomes is quantified by the Fisher information (FI) per detected photon
[…]
which vanishes in the limit $\epsilon\to 0$ [14]. The precision, quantified by the variance of an unbiased estimator $\hat{\epsilon}$ is lower-bounded by the Cramér–Rao lower bound (CRLB) $\text{Var}(\hat{\epsilon})\geq 1/[N\mathcal{F}(\epsilon)]$, where $N$ represents the number of detected photons. The estimator variance goes to infinity in the limit $\epsilon\to 0$, which suggests that resolving arbitrarily small separations becomes infeasible with DI measurements.
An optimal measurement, in contrast, employs the coherent temporal mode filtering that project the input signal onto an orthonormal set of HG modes. This task is analogous to the two-point resolution problem in spatial imaging. For a Gaussian point spread function, this mode basis has been shown to saturate the quantum FI - the ultimate limit of precision attainable over all possible measurements for a given input quantum state - in the context of separation estimation [14]. Projecting the input signal onto the HG basis in the frequency domain is equivalent to filtering its temporal profile - the Fourier transform of its spectrum - with the corresponding temporal HG basis. The probability of detection in mode $n$ is
$P(n|\epsilon)=\frac{\epsilon^{2n}}{16^{n}n!}\exp\left(-\frac{\epsilon^{2}}{16}\right).$ (1)
And the FI for this HG measurement basis can be calculated as
$\mathcal{F}_{\text{HG}}(\epsilon)=\sum_{n=0}^{\infty}\frac{1}{P(n|\epsilon)}\left(\frac{\partial P(n|\epsilon)}{\partial\epsilon}\right)^{2}\approx\frac{1}{4}.$ (2)
The constant FI saturates the quantum FI for all $\epsilon$, verifying the optimality of the HG measurement basis and enabling super-resolution.
As the HG measurement is particularly advantageous in small separations, we will focus on the precision and ability to resolve small separations under limited photon budgets in realistic experiments. For small separations, the projection onto the $\text{HG}_{0}$ and $\text{HG}_{1}$ mode contributes most of the FI. According to Eq. 1, the ideal projection probabilities onto HG0 and HG1 modes yield $P(1|\epsilon)/P(0|\epsilon)=\epsilon^{2}/16$. This ratio allows for the construction of a simple raw estimator $\hat{\epsilon}_{\text{raw}}=4\sqrt{N_{1}/N_{0}}$, where $N_{0}$ and $N_{1}$ are the experimentally measured counts of HG0 and HG1 projections, respectively.
In practical implementations, the measured projection probabilities will inevitably deviate from the ideal values predicted by Eq. 1 - a discrepancy that can be interpreted as mode crosstalk. To model this, we account for mode crosstalk between HG0 and HG1 measurements via the coupling matrix
$\bm{M}=\begin{pmatrix}\alpha&1-\beta\\ 1-\alpha&\beta\end{pmatrix}$ (3)
where $\alpha,\beta\in[0,1]$ quantify the mode filtering fidelity. The theoretical probabilities $\bm{p}=(P(0|\epsilon),P(1|\epsilon))^{\top}$ are thus perturbed to $\bm{\tilde{p}}=\bm{Mp}$ in the measurement basis, after which $\bm{\tilde{p}}=(\tilde{P}(0|\epsilon),\tilde{P}(1|\epsilon))^{\top}$ is normalized. Assuming low mode crosstalk $\alpha\approx 1$ and $\beta\approx 1$, we derive the FI for small separations using Eq. 2 as (see the Supplementary Materials for derivations)
$\mathcal{F}(\epsilon)\approx\frac{1}{4\left[(1-\alpha)/(\epsilon/4)^{2}+1\right]}.$ (4)
It can be seen from Eq. 4 that the FI in the presence of mode crosstalk is critically dependent on the leakage from the HG0 mode into the HG1 mode. If $(\epsilon/4)^{2}\ll 1-\alpha$, the FI decreases significantly from the ideal case. Low-crosstalk mode filtering is therefore crucial for resolving very small separations. In the presence of crosstalk, although $\mathcal{F}(\epsilon)\rightarrow 0$ as $\epsilon\rightarrow 0$, Eq. 4 still gives an advantage in FI over DI methods when the crosstalk is small.
To estimate $\epsilon$ in the presence of crosstalk, we employ maximum likelihood estimation (MLE) using the crosstalk-perturbed probabilities $\tilde{P}(0|\epsilon)$ and $\tilde{P}(1|\epsilon)$. The MLE finds the parameter value that maximizes the likelihood function $\mathcal{L}(\bm{x}|\epsilon)=\prod_{i=1}^{N}\tilde{P}(x_{i}|\epsilon)$ for observing the measured data $\bm{x}=\{x_{1},x_{2},\ldots,x_{N}\}$. The estimation is performed subject to the parameter space constraint $\Theta=\{\epsilon\in\mathbb{R}\mid\epsilon\geq 0\}$. For small separations, the MLE estimators can be written as
$\hat{\epsilon}_{\text{MLE}}=\operatorname*{arg\,max}_{\epsilon}[N_{0}\ln{\tilde{P}(0|\epsilon)}+N_{1}\ln{\tilde{P}(1|\epsilon)}].$ (5)
The MLE is asymptotically unbiased and efficient in the limit $N\to\infty$, implying its variance approaches the CRLB.
In real experiments with finite statistics, the MLE for this separation estimation problem exhibits a non-zero bias, particularly at small separations. This bias comprises contributions from both higher-order asymptotic terms [53] and a distinct positive bias arising from the parameter space constraint. In our model, the separation $\epsilon$ is defined in the non-negative domain, $\Theta$. However, due to shot noise, the experimentally observed normalized counts $\bm{f}=(N_{0}/N,N_{1}/N)^{\top}$ may lie outside the physically valid range predicted by our model $\bm{\tilde{p}}$. When this occurs, the likelihood function $\mathcal{L}(\bm{x}|\epsilon)$ takes its maximum at the boundary of the parameter space constraint (i.e. $\epsilon=0$), leading to a positively skewed distribution of the estimator. This skewing effect results in an unavoidable bias for $\epsilon\to 0$, with the magnitude of the bias determined by the estimator’s distribution. The bias diminishes with higher photon counts $N$, as statistical fluctuations on $N_{0}$ and $N_{1}$ decrease, making the estimator’s distribution at the boundary less probable (see the Supplementary Materials for details).
To assess our measurement apparatus, we quantify the estimation performance by the mean squared error (MSE) of the estimator, $\text{MSE}(\epsilon,N)=\braket{(\hat{\epsilon}-\epsilon)^{2}}$. Importantly, the MSE incorporates both the variance and the squared bias of the estimator, $\text{MSE}(\epsilon,N)=\text{Var}(\epsilon)+b(\epsilon,N)^{2}$, being lower bounded by the CRLB with a bias term [54]
$\text{MSE}(\epsilon,N)\geq\frac{[1+b^{\prime}(\epsilon,N)]^{2}}{N\mathcal{F}(\epsilon)}+b(\epsilon,N)^{2},$ (6)
where $b^{\prime}(\epsilon,N)$ represents the derivative of $b(\epsilon)$ with respect to $\epsilon$. Only as $b(\epsilon)\rightarrow 0$ does the MSE converge to the standard unbiased CRLB, $\text{Var}(\hat{\epsilon})\geq 1/[N\mathcal{F}(\epsilon)]$. From the MSE, we quantify the sensitivity of the apparatus as the minimal resolvable separation $\epsilon_{\text{min}}$ for which the parameter-to-error ratio (PER) of the estimate,
$\text{PER}(\epsilon,N)=\epsilon^{2}/\text{MSE}(\epsilon,N),$ (7)
is greater than 1 [22]. This minimal resolvable separation defines a modified Rayleigh criterion for resolving spectral lines under certain detected photon budgets. In this work, we employ MSE and PER as the figures of merit to quantify the performance of our experimental platform.
II.2 Mode-selective Raman quantum memory
[…]
In the low-coupling regime (i.e. at low storage efficiencies), the memory operates in a single-mode fashion: the temporal mode of the signal stored in the memory is directly determined by the temporal mode of the control pulse [56] (see the Supplementary Materials for more discussions). As illustrated in Fig.1(b), when both the signal and control fields are in the HG0 mode (labeled as 1), the signal is stored and subsequently retrieved. In contrast, when the signal is in the HG0 mode but the control is in the HG1 mode (labeled as 2), there is nearly no storage or retrieval. Experimental results for these two scenarios are presented in Fig.1(c). To quantify the mode selectivity, we measured the total efficiencies for all combinations of signal and control pulses prepared in HG0 and HG1 modes. The efficiencies were normalized for each signal mode, producing the mode crosstalk matrix shown in Fig. 1(d). Our results confirm high mode selectivity: we achieved $99.61\pm 0.48\%$ mode selectivity for the storage of HG0 signal with HG0 control, and a selectivity of $0.39\pm 0.02\%$ when stored with HG1 control. These results were obtained using both read-in and read-out control pulses with energies of $125\pm 1$ pJ, yielding total efficiencies of approximately 11% along the diagonal terms. It is important to note that at higher coupling strengths (e.g., with stronger control pulses), this single-mode behavior degrades due to the increased storage of temporal modes orthogonal to that of the intended control field. This highlights an inherent trade-off between mode selectivity and overall memory efficiency. By achieving low crosstalk and high mode selectivity in the low-coupling regime, our Raman memory provides a robust and effective platform for the super-resolution task.
II.3 Experimental superresolution with mode-selective quantum memory
Figure 2: Experimental estimation results. This graph compares the maximum likelihood estimators (red markers) with raw estimators (yellow markers), both derived from $10\times 10^{3}$ total detected photons. Error bars represent one standard deviation from 50 bootstrapped samples. The dashed black line indicates the ground truth separation. The shaded regions illustrate the theoretically predicted standard deviations from the quantum CRLB (blue) and the conventional direct intensity detection CRLB (gray).
To perform estimation experiments using the mode-selective Raman quantum memory described above, we first prepared the signal as an incoherent mixture of two Gaussian spectral lines, each with a linewidth of $\sigma=5.30$ MHz (see Methods for detailed procedure). We varied their frequency separation across 21 values, ranging from 0 to 1 in steps of 0.05. For each signal, we performed two measurements, storing the signal with an HG0 or an HG1 control read-in pulse. In both cases, retrieval was conducted using a $\text{HG}_{1}$ control pulse after 200 ns storage and the retrieved signal was detected with superconducting nanowire single-photon detectors. At $\epsilon=0$, the storage efficiency and total internal efficiency for the $\text{HG}_{0}$ control read-in were measured to be $39.5\pm 0.6\%$ and $14.6\pm 0.1\%$ respectively, using control pulses with energies of $130\pm 1$ pJ. For each separation, we recorded approximately $N\approx 2\times 10^{5}$ detector clicks in total. To estimate the frequency separation $\epsilon$, we first performed a full calibration of the experimental system with $1.6\times 10^{5}$ detected counts for each separation. By fitting the experimental model to this calibration data, we determined the measured crosstalk from the $\text{HG}_{0}$ to $\text{HG}_{1}$ modes in the estimation experiment to be 0.34%. It is noteworthy that this is a separate experiment from the demonstration of mode filtering described in the previous section, accounting for the slight difference in the measured mode crosstalk values. With the calibrated model, in each run we performed MLE to compute the estimate $\hat{\epsilon}_{\text{MLE}}$ based on Eq. 5 with registered counts $N_{0}$ and $N_{1}$. We then obtained the MSE for each separation by averaging $(\hat{\epsilon}_{\text{MLE}}-\epsilon)^{2}$ over 50 bootstrapped experiments with the same $N$ from the total dataset of detector clicks. The uncertainty in the MSE was calculated by repeating this procedure 10 times. The detailed procedure of data collection and analysis is described in Methods.
The estimation results, including estimator standard deviations, for a total of $N=10\times 10^{3}$ detected counts are shown in Fig. 2 (the results for $2\times 10^{3}$ and $100\times 10^{3}$ counts are provided in the Supplementary Materials). The yellow markers represent the raw estimators $4\sqrt{N_{1}/N_{0}}$, which exhibit a systematic bias primarily due to the experimental mode crosstalk arising from the memory storage process and control field leakage. While these biased estimators possess low variances, their systematic bias compromises their accuracy. In contrast, the MLE estimators exhibit much lower bias compared to the raw estimator. For the smallest separations, such as $\epsilon=0$ and $0.05$, the MLE shows a non-zero bias with increased variances. The relatively larger variances (reflected by the error bars) for these small-separation MLE estimates, compared to the raw estimators, clearly demonstrate the trade-off between estimator variance (precision) and bias (accuracy), as elucidated in Eq. 6. Importantly, despite being above the ideal quantum bound, all MLE error bars consistently lie within the DI bound, providing a clear demonstration of enhanced precision relative to DI methods. To validate the platform’s unique functionality of on-demand storage, retrieval and coherent mode conversion, we also performed experiments and demonstrated super-resolving estimation with storage durations from 150 ns to 250 ns and using HG0 and HG1 retrieval modes. The experimental results are given in the Supplementary Materials.
Figure 3: Estimation performance under different photon budgets. (a) Experimental MSE times photon number $N$. The plot shows $\text{MSE}\times N$ for total detected photon counts of $2\times 10^{3}$ (orange triangles), $10\times 10^{3}$ (green squares), and $100\times 10^{3}$ (blue circles). MSE error bars are determined from 10 bootstrapped results. The solid red curve represents the theoretical CRLB for mode filtering using our Raman quantum memory, incorporating measured mode crosstalk. As the photon number increases, the MSE approaches the CRLB line. For comparison, the quantum CRLB (blue dash-dotted line) and the direct intensity detection CRLB (gray dashed line) are also shown. (b) Parameter-to-error ratio for different total detection counts. Solid lines represent theoretical PER values calculated using Eq. 7, derived from theoretical MSE values that incorporate the measured crosstalk parameters. The corresponding symbols (circles, squares, and triangles) indicate PER values computed from our experimentally measured MSE. The sensitivity, or smallest resolvable separation, is defined as the point where the PER exceeds 0 dB. Higher detection counts consistently lead to improved PER and enable the resolution of smaller frequency separations. Specifically, for $N=100\times 10^{3}$ photons, our system achieves a smallest resolvable separation of $\epsilon=0.05$.
The residual bias observed for near-zero separations in MLE is directly attributed to the non-negativity constraint imposed on the parameter $\epsilon$, as we analyzed in Section A. For the limited photon budgets in our experiments ($2\times 10^{3}$ to $100\times 10^{3}$), this estimator bias is unavoidable for small separations and is dependent on the number of detected photons. Figure 3(a) illustrates the measured MSE and the CRLB, both scaled by $N$ for comparison across different total detection counts. For separations $\epsilon<0.5$, the MSE is significantly lower than the CRLB for the DI method. The experimentally measured MSE quantifies the total estimation error for a given photon budget, including contributions from both the bias and the variance of the estimator. At low photon numbers, such as $N$ = $2\times 10^{3}$, the MSE falls below the CRLB for small separations $\epsilon<0.15$ — a consequence of the estimator bias. As the photon count increases to $100\times 10^{3}$, the MSE approaches the CRLB as the bias decreases. This observation is consistent with our earlier discussion of MLE bias at small separations. These results show that for moderate photon budgets, the MLE exhibits a non-zero bias for small separations. However, the contribution of the bias to the MSE becomes negligible with increasing photon counts, allowing the estimator to perform reliable estimation at higher photon numbers.
II.4 Sensitivity and precision enhancement benchmarking
Figure 4: Precision enhancement benchmarking. (a) Precision enhancement as a function of spectral line separation $\epsilon$. The red curve shows the theoretical precision enhancement, calculated as the ratio of the mode filtering CRLB to the direct intensity measurement CRLB. The symbols represent experimentally measured precision enhancements based on the MSE values obtained at various total detection counts: $N=2\times 10^{3}$ (orange triangles), $10\times 10^{3}$ (green squares), and $100\times 10^{3}$ (blue circles). The average of data points lie slightly below the theoretical curve due to additional MSE arising from experimental fluctuations. The superresolution parameter $\mathfrak{s}$ is defined as the asymptotic limit of the red curve as $\epsilon\rightarrow 0$. (b) Comparison of superesolution parameters for various schemes. This panel benchmarks the performance of our system against other super-resolution techniques and DI methods. QPG: quantum pulse gate [40]; PuDTAI: pulsedivision time-axis-inversion [43]; SUSI: super-resolution via spectral inversion [45]; QMTI: quantum-memory temporal imaging [43]; FT: Fourier transform spectrometer (Bruker IFS 125HR); Grating: grating spectrometer of grating length 1 cm and grating period of 1200 mm.
To benchmark the sensitivity of our measurement scheme, we examine the minimum resolvable separation based on the PER ratio, as defined in Eq. 7. Figure 3(b) shows the PER as a function of separation for different total detection counts. Our experimental results demonstrate a PER$=4.4\pm 0.5$ dB for $\epsilon=0.05$ with $100\times 10^{3}$ photon counts. This signifies that our system can resolve spectral line separations as small as 265 kHz for spectral lines with a linewidth of 5.30 MHz. In contrast, for lower photon numbers ($N=2\times 10^{3}$ and $10\times 10^{3}$), the PER remains below 0 dB at $\epsilon=0.05$, implying that the frequency separation cannot be reliably distinguished from a zero separation at these photon levels. The solid lines in Fig. 3(b) represent the theoretical PER values, derived from theoretical MSE simulations, which show good agreement with the experimentally measured PER.
In addition to sensitivity, we evaluate the precision enhancement enabled by our memory-based measurement apparatus with mode filtering, benchmarking its performance against DI methods. We quantify the enhancement as the ratio of the FI for the two methods in the limit of vanishing separation $\epsilon\to 0$. This limiting value corresponds to the superresolution parameter, $\mathfrak{s}$, which is introduced by Mazelanik et al. [43] as
$\mathfrak{s}=\lim\limits_{\epsilon\rightarrow 0}(\mathcal{F}/\mathcal{F}_{\text{DI}}).$
This parameter defines the ultimate limit of precision enhancement achievable with an infinite photon budget, as a truly unbiased estimator is only realized when $N\to\infty$. This parameter can be used to assess and compare the performance of various super-resolving measurement apparatuses, with its value primarily determined by mode crosstalk. The red curve in Fig. 4(a) illustrates how the theoretical precision enhancement, represented by $\mathcal{F}/\mathcal{F}_{\text{DI}}$, varies with frequency separation for our platform. In the limit $\epsilon\rightarrow 0$, the ratio asymptotically approaches a value of approximately 37. Using this definition, our system achieves the highest reported superesolution parameter value for frequency separation estimation to date [40, 43, 45], as shown in the comparative graph in Fig. 4(b). For comparison, we also show the performance of DI methods, including quantum-memory temporal imaging [43], a Fourier transform spectrometer (Bruker IFS 125HR) and a grating spectrometer; all of these fall below $\mathfrak{s}=1$. We should also highlight that while the super-resolving methods presented span various bandwidth regimes, our Raman memory uniquely operates in the MHz to GHz range, a bandwidth dictated by the hyperfine splitting between the initial and storage states and the memory storage time.
It is important to note, however, that this theoretical precision enhancement $\mathfrak{s}$ in the $\epsilon\rightarrow 0$ limit is experimentally unattainable, as achieving unbiased estimation in this regime requires an infinite number of detected photons. In practice, when the estimator exhibits negligible bias, the enhancement can be estimated using the ratio of the measured mode filtering MSE to the CRLB of DI methods. In Fig. 4(a), we plot these ratios as data points at different total photon counts. These data points align well with the tendency of the theoretical prediction, though subject to experimental imperfections and fluctuations. However, at very small separations, the estimator bias can cause the measured MSE to fall below the CRLB, leading to calculated enhancement ratios that may exceed the theoretical values (red curve in Fig. 4(a)). In this region, we expect both the mode filtering method and the DI detection method exhibit estimator biases, making the determination of realistic precision enhancement more complex [21]; a detailed analysis will be explored in future work. Our experimental measurements demonstrate a practical $(34\pm 4)$-fold improvement at $\epsilon=0.05$ and a $(28\pm 6)$-fold improvement at $\epsilon=0.1$ using $100\times 10^{3}$ detected photons, confirming the feasibility and effectiveness of the method. Furthermore, we expect DI measurements, which inherently have larger estimator variances from the FI analysis, require higher number of photons to achieve nearly unbiased estimation compared to the optimal mode filtering. The mode filtering method, with its lower variance, is especially valuable for photon-limited applications where conventional techniques fail to achieve accurate estimation.
III Discussion
The experimental results demonstrate a clear improvement in the precision of estimating spectral line separation. Our system achieves a theoretical 37-fold precision enhancement in the limit $\epsilon\rightarrow 0$, and a practical $(34\pm 4)$-fold enhancement at $\epsilon=0.05$, relative to DI method for sub-linewidth separations. Our system exhibits sensitivity sufficient to resolve frequency separations as small as 1/20 of the individual signal linewidth. Among TF superresolution techniques [40, 43, 45], our Raman memory operates over a unique bandwidth, from MHz to GHz, where no classical strategy is available. Going beyond resolving two spectral lines, our platform’s fully programmable temporal mode filtering also supports the capability of resolving more complex spectral features. The user-defined mode filtering is pivotal for multi-parameter estimation, where the optimal measurement basis can be complicated depending on the task [57, 58, 59, 60]. Such optimal measurement bases are ideally realized through TF mode sorting onto multiple temporal modes [42], which could be implemented using multiple cascaded memories or a loop architecture with a single memory.
As shown in Eq. 4, achieving even higher precision in our system hinges on reducing mode crosstalk, which primarily stems from the memory storage process. Crucially, there is a trade-off between mode crosstalk and storage efficiency, which arises from the low-coupling requirement for the memory’s single-mode behavior (see the Supplementary Materials for simulation results). To address this limitation, our recently developed efficiency enhancement via light-matter interference (EEVI) protocol [61] offers a path to realizing high-efficiency storage and retrieval while preserving single-modeness. Similarly, cavity-enhanced Raman memories [62, 63] can offer similar efficiencies, though their bandwidth is constrained by cavity finesse. Furthermore, employing optimal control techniques [64, 65] to tailor control pulse shapes can maximize storage efficiency for the target mode while simultaneously reducing it for orthogonal modes, thereby minimizing crosstalk.
Despite the importance of precision enhancement, the practical advantage of the super-resolving measurement, quantified by FI per input photon, is significantly impacted by photon loss of the setup. To date, all experimental platforms for TF super-resolution suffer from low end-to-end efficiency. In our current setup, we measure an overall efficiency of approximately 0.3%, despite achieving an internal total efficiency of $14.6\pm 0.1\%$. This disparity primarily arises from the stringent filtering needed to suppress the co-propagating control field, which is detuned by just 9.2 GHz from the signal. To achieve an extinction ratio of $\sim 4\times 10^{7}$, we used a calcite beam displacer, a Glan-Taylor polarizer, and three double-passed Fabry–Pérot etalons, each contributing to the overall photon loss. Notably, only changes in the internal efficiency alter mode crosstalk. In contrast to Raman memories, ladder-type quantum memories, such as fast-ladder memory [66] or off-resonant cascaded absorption memory (ORCA) [67, 68], exploit a higher-lying excited state and widely separated, counter-propagating signal and control fields. This configuration, allowing efficient suppression of control leakage without complex filtering, has demonstrated noise per signal photon as low as $10^{-5}$ [67] with end-to-end efficiencies up to 35% [66]. Integrating our high-fidelity mode filtering with these architectures holds the potential for realizing unconditional TF super-resolution.
In conclusion, we have demonstrated a frequency superresolution protocol based on mode-selective Raman quantum memory. The technology developed here holds promise for diverse applications, including high-precision clock synchronization [69], spacetime positioning [70, 71], photon dose-limited metrological tasks [7], and processing of TF encoded quantum information [6, 28, 27, 29]. Future works will focus on further improving precision through the implementation of the EEVI protocol and control optimization techniques. We will also explore the potential of ORCA memories to achieve unconditional super-resolving measurements. Furthermore, we aim to demonstrate the platform’s capabilities for multi-parameter estimation and its integration into quantum sensor networks. The unique combination of high-fidelity mode filtering, on-demand storage and mode conversion capabilities positions our platform as a multifunctional tool for advanced TF metrology, paving the way for integrated quantum sensors tailored to photon-limited applications and distributed sensing networks.
IV Methods
Experimental setup
Figure 5: (a) Experimental setup. The apparatus consists of three main sections: pulse generation, memory interaction, and filtering and detection. Both signal and control pulses originate from a continuous wave (CW) laser at the control-field frequency. An electro-optic modulator (EOM), EOM 1, first carves the CW light into pulses. A 90:10 beam splitter (BS) then splits the laser into a high-power control path (90%) and a low-power signal path (10%). The laser in the signal path is later shifted to the signal frequency via EOM 2. The two beams are recombined via a beam displacer (BD) before entering the memory cell for storage and retrieval. Post-retrieval filtering comprises a Glan-Taylor polarizer (GTP), three double-passed etalons, and a band-pass filter (BP). The signal is finally routed to fiber beam-splitters and four superconducting nanowire single-photon detectors (SNSPDs) for detection. Employing multiple detectors helps mitigate detector dead time effects and enables higher overall count rates. (AWG: arbitrary waveform generator; PC: Pockels cell; PBS: polarizing beamsplitter; M: mirror; DM: dichroic mirror; HWP: half-wave plate; WPs: a half-wave plate and a quarter-wave plate.) (b) Pulse sequence. The sequence consists of three pulses: the first pulse is the signal input, which carries the spectral information to be estimated. The second pulse serves as the control read-in pulse, initiating the coherent storage process. The third pulse acts dually as both the control read-out pulse to retrieve the stored signal and a reference pulse for real-time signal power monitoring. The signal sequence is delayed by 437 ns to ensure temporal overlap of the signal input pulse with the control read-in pulse at the memory cell for storage. After a 200 ns storage period, the stored signal was retrieved using the control read-out pulse. During the measurement window, the pump laser was disabled, while the Pockels cell was activated exclusively during the retrieval window to suppress background from the pre-storage signal.
Our Raman memory setup is shown in Fig. 5(a). The storage medium consists of a 75 mm-long cesium (Cs-133) vapor cell with a three-layer $\mu$-metal magnetic shielding. We heated the cell to $105\pm 1$°C using a quad-twisted cryogenic wire, achieving an effective optical depth of $(4.8\pm 0.1)\times 10^{3}$. The atoms were initialized to the initial state |$\ket{g}$ by an external cavity diode laser (ECDL, Toptica DL100) driving the D1 transition ($6^{2}\textbf{S}_{1/2}\rightarrow 6^{2}\textbf{P}_{1/2}$). An electro-optical modulator (EOM 3, EOSpace 850UL) gated the pump laser, switching it off during the experimental window. We achieved a pumping efficiency of $(99.6\pm 0.1)$%.
We operated our memory with a detuning $\Delta=2\Delta_{\text{hf}}$ to suppress four-wave-mixing noise [55]. Both the signal and control fields were derived from a single continuous wave (CW) ECDL laser (Toptica DL Pro) operating at the control frequency (351.7122 THz). An arbitrary waveform generator (AWG, Tektronix AWG70001A) drove the electro-optical modulator EOM 1 (Sacher Lasertechnik AM830PF) to carve the CW laser into pulses. The laser was then split using a 90:10 beam splitter, directing 90% of the optical power to the control field path. The remaining 10% went to EOM 2 (New Focus 4851), which generated a sideband at the signal frequency (351.7030 THz). We selected this sideband as the signal field by using a Fabry-Pérot etalon with a free spectral range (FSR) of 36.8 GHz. The signal field was delayed by 430 ns using a long fiber to achieve temporal synchronization with the control pulses, ensuring their temporal overlap for memory storage and retrieval. Before entering the memory, the signal field also passed through a Pockels cell to remove residual CW background in the retrieval window. The polarization of the signal and control fields was set to orthogonal for memory interaction, allowing their combination using a beam displacer before being focused into the vapor cell.
We used a pair of convex lenses to focus the signal, control and pumping beams into the vapor cell to enhance the interaction strength. The beam widths at the focus were approximately $164\pm 5$ $\mu$m for the signal and $190\pm 5$ $\mu$m for the control. As shown in Fig. 5(b), the first signal pulse meets the second control pulse for storage. After a storage time of 200 ns, the second control pulse retrieves the stored signal. The third signal pulse serves as a reference for the signal pulse power. All signal and control pulses share the same Gaussian envelope width parameter as a fundamental Gaussian pulse (HG0) with an electric field full width at half maximum (FWHM) of 50 ns.
After the memory, we suppressed the control and pump fields using a Glan-Taylor polarizer, followed by three double-passed etalons (two with FSR of 18.4 GHz and one with FSR 103 GHz) and a bandpass filter (central wavelength: 850 nm, FWHM Bandwidth: 10 nm). Finally, we split the signal into four equal parts and detected them using four superconducting nanowire single-photon detectors (SNSPDs, Photon Spot). The detected counts were registered using a time tagger (Swabian TimeTagger20).
Signal preparation
The signal in our separation estimation task is an incoherent mixture of two Gaussian spectral lines. Its frequency-domain representation can be written as
$\Psi(\omega)=\frac{1}{\sqrt{2}}$ $\left[\exp\left({-i\frac{\phi}{2}}\right)\right.\psi(\omega-\omega_{0}-\frac{\epsilon\sigma}{2})$ (8)
$+\left.\exp\left({i\frac{\phi}{2}}\right)\psi(\omega-\omega_{0}+\frac{\epsilon\sigma}{2})\right],$
where $\phi\in(-\pi,\pi]$ is a random phase between the two spectral lines. Here, we carved the temporal profile of our signal field as the inverse Fourier transform of Eq. 8,
$\Psi(t)=A\cos\left(\frac{\epsilon\sigma t-\phi}{2}\right)\exp(-t^{2}\sigma^{2}),$
where $A$ is the amplitude. The signal pulse has a spectral width of $\sigma=33.3$ Mrad/s (5.30 MHz). In our experiment, we generated signals with separation parameters $\epsilon$ from 0 to 1 with an increment of 0.05 (0.265 MHz). To introduce the incoherence of the two spectral lines, we prepared the signal in a mixture of four different relative phases, $\phi=-\pi/2,0,\pi/2,\pi$, for each separation.
One key requirement for this experiment is the accurate generation of signal and control pulses. To ensure high-fidelity pulse carving, we characterized the frequency response of our electronic pulse generation system, which comprises the AWG, an RF splitter, and two RF amplifiers that drive the EOM 1. By measuring the system’s frequency response function, we applied a correction to the input signals of the AWG, compensating for frequency-dependent variations in the RF components. This correction ensures uniform amplification across all frequency components, minimizing distortions in the carved pulses. The measured intensity pulse shapes of HG0 and HG1 are presented in the Supplementary Materials.
Data collection and analysis
For each signal, comprising one separation and one phase setting, we first stored it using a control pulse of HG0 with a pulse energy of $130\pm 1$ pJ. At this control energy, our memory typically operates with a storage efficiency of $(39.5\pm 0.6)\%$ and a total internal efficiency of $(14.6\pm 0.1)\%$ for a storage duration of 200 ns. The stored signal was retrieved using a control pulse in the HG1 mode, with the same pulse energy as the HG0 write-in pulse. The HG1 mode was chosen to minimize distortion in the pulse sequence due to its symmetrical temporal profile. Importantly, the choice of retrieval mode has a negligible impact on the retrieval efficiency. The pulse sequence was repeated every 3 $\mu$s, with a total measurement time of 2 s. To accurately account for signal background noise and dark counts, we also recorded data with the control pulse blocked for an equivalent duration. The counts detected in the retrieval time window during this control-blocked run were later used to subtract the noise counts from the retrieved signal counts to minimize the estimator bias caused by these noises. The above procedures were repeated for all four phases. Subsequently, we modified the control read-in pulse to the HG1 mode while keeping the read-out pulse unchanged and repeated the measurements. Finally, the entire experiment was repeated for all separations of interest.
To obtain the final detected counts, we first identified a reference pulse count that yielded approximately $N$ = $2\times 10^{3}$, $10\times 10^{3}$, or $100\times 10^{3}$ total retrieval counts (with a variation less than 5%) of the signals across all combinations of the two control read-in pulses and four phases. For each separation, we used the reference counts to randomly sample a subset that yielded $N$ retrieved counts from the data files. The corresponding noise counts were subtracted from the retrieved counts to correct the background contributions. This random sampling approach, combined with the use of reference counts, was employed to mitigate the impact of signal power fluctuations on the final estimation process. The retrieval counts acquired for HG0 and HG1 storage, $N_{0}$ and $N_{1}$ respectively, were used directly to compute the raw estimator. We repeated the entire sampling and analysis procedure 50 times using bootstrapping to estimate the variances and MSEs of the estimators. To obtain the error bars for the MSE shown in Fig. 3, we repeated the bootstrapping process 10 times to get 10 datasets and calculated the standard deviations of the MSE.
System calibration
To characterize the crosstalk matrix $\bm{M}$ of the experimental system and the perturbed HG projection probabilities $\tilde{P}(0|\epsilon)$ and $\tilde{P}(1|\epsilon)$ in Eq. 5, we used a dataset for separations $\epsilon$ from 0 to 1 with $N=1.6\times 10^{5}$ detected photon counts for each separation. For each separation, we normalized the retrieved counts $N_{0}$ and $N_{1}$ corresponding to the two HG projections to obtain the relative frequencies $f_{0}=N_{0}/(N_{0}+N_{1})$ and $f_{1}=N_{1}/(N_{0}+N_{1})$. The values of $\alpha$ and $\beta$ were estimated by minimizing the least square cost
$\begin{split}C(&f_{0},f_{1}|\alpha,\beta)\\ &=\sum_{\epsilon}\left\{\left[f_{0}(\epsilon)-\tilde{P}(0|\epsilon)\right]^{2}+\left[f_{1}(\epsilon)-\tilde{P}(1|\epsilon)\right]^{2}\right\}.\end{split}$
The perturbed probabilities, parameterized by $\alpha$ and $\beta$, were compared to the measured relative frequencies, and the sum of squared residuals across all separations was computed. We then optimized the parameters using nonlinear least-squares minimization. The resulting values characterize the performance of our experimental system and were subsequently used to compute the MLE estimates.
Acknowledgments
We acknowledge J. Szuniewicz for help in developing the EOM control code, R. B. Patel for help in setting up the detectors. We thank B. Brecht, M. G. Raymer, A.M. Steinberg and S. Yu for helpful discussions. This work was supported by the European Union’s Horizon 2020 Research and Innovation Programme Grant No. 899587 Stormytune and the Engineering and Physical Sciences Research Council via the Quantum Computing and Simulation Hub (Grant No. T001062). A. Z. acknowledges a UK Research and Innovation Guarantee Postdoctoral Fellowship under the UK government’s Horizon Europe funding Guarantee (EP/Y029127/1). S.E.T. acknowledges an Imperial College Research Fellowship.
Supplementary Material: Super-resolving frequency measurement with mode-selective quantum memory
IV.1 Fundamental bias at small separations
In the maximum likelihood estimation process, the non-negativity constraint of the parameter space $\Theta$ introduces a positive bias at very small frequency separations, where the distribution of an ideally unbiased estimator may extend into the negative region. Here, we present both an analytical analysis and numerical simulations to characterize this fundamental bias of a two-mode demultiplexing system with crosstalk matrix $\bm{M}$ (Eq. 3 in the main text).
With the crosstalk matrix $\bm{M}$, the normalized measurement probability of HG1 projection is
$p_{1}=\beta-\frac{16(\alpha+\beta-1)}{16+\epsilon^{2}},$ (S1)
and the HG0 projection probability is $p_{0}=1-p_{1}$. The exact expression for the Fisher information of the non-ideal two-mode projection method can be derived from Eq. 2 in the main text as follows:
$\mathcal{F}(\epsilon)=\frac{(\alpha+\beta-1)^{2}(\epsilon/4)^{2}}{4(1+(\epsilon/4)^{2})^{2}(\alpha+(1-\beta)(\epsilon/4)^{2})(1-\alpha+\beta(\epsilon/4)^{2})}.$ (S2)
We assume the mode crosstalk between the two modes is low, i.e. $\alpha\approx 1$ and $\beta\approx 1$, and the separation is small. Then Eq. S2 can be approximated as
$\mathcal{F}(\epsilon)\approx\frac{1}{4\left[(1-\alpha)/(\epsilon/4)^{2}+1\right]}.$ (S3)
Using maximum likelihood estimation (Eq. 5 in the main text), the likelihood function is given by
$\mathcal{L}=N_{0}\ln{p_{0}}+N_{1}\ln{p_{1}},$ (S4)
where $N_{0}$ and $N_{1}$ represent the experimentally measured detection counts for HG0 and HG1 projections, respectively, with the total count given by $N=N_{0}+N_{1}$. The analytical solution of the MLE estimator can be obtained by solving the critical point equation $d\mathcal{L}/d\epsilon=0$, yielding $\hat{\epsilon}=\sqrt{\frac{16(N_{1}-(1-\alpha)N)}{\beta N-N_{1}}}$. However, this solution is valid only when both the numerator and denominator are non-negative; thus, we require $N_{1}\geq(1-\alpha)N$ (We neglect the condition $N_{1}<\beta N$, as the probability of $N_{1}>\beta N$ is negligible for small separations and low mode crosstalk). This condition corresponds to the range of the modal prediction discussed in the main paper. If a combination of $N_{0}$ and $N_{1}$ values falls outside this range, the likelihood function $\mathcal{L}$ achieves its maximum at the non-differentiable boundary point $\epsilon=0$. In such cases, the MLE defaults to $\hat{\epsilon}=0$. Therefore, the MLE estimators are
$\hat{\epsilon}_{\text{MLE}}=\begin{cases}\sqrt{\frac{16(N_{1}-(1-\alpha)N)}{\beta N-N_{1}}},&\text{if }N_{1}\geq(1-\alpha)N,\\ 0,&\text{if }N_{1}<(1-\alpha)N.\end{cases}$ (S5)
The maximum likelihood estimation method inherently produces only non-negative estimators based on the detected counts. When $N_{0}$ and $N_{1}$ result in a negative estimate, the estimator is set to zero. Consequently, this introduces a bias toward positive values, as all negative estimates are truncated to zero.
Figure S1: Simulation results of the mean square error (MSE) and the unbiased Cramér-Rao lower bound (CRLB) for maximum likelihood estimation. (a) Simulated MSE, where the white region indicates the onset of truncation due to the non-negativity constraint. (b) MSE with the non-negativity constraint applied. (c) The unbiased CRLB. (d) The difference between the CRLB and the MSE (CRLB -- MSE).
Since we employ photon count detection, the detection counts $N_{0}$ and $N_{1}$ follow a Poisson distribution. Specifically, we model $N_{1}$ as $\text{Poisson}(\mu_{1})$, where $\mu_{1}=p_{1}N$. Substituting this distribution into the estimator formula in Eq. S5, we derive the mean square (MSE) error as
$\text{MSE}(\epsilon,N)=\sum_{k=0}^{(1-\alpha)N-1}\epsilon^{2}\cdot\frac{\mu_{1}^{k}e^{-\mu_{1}}}{k!}+\sum_{k=(1-\alpha)N}^{\beta N-1}\left(4\sqrt{\frac{k-(1-\alpha)N}{\beta N-k}}-\epsilon\right)^{2}\cdot\frac{\mu_{1}^{k}e^{-\mu_{1}}}{k!}.$ (S6)
(Here, we neglected the $N_{1}>\beta N$ part, as it is extremely unlikely to occur for small separations when $\beta\approx 1$.)
The MSE simulated using our experimentally measured crosstalk is shown in Fig. S1 as a function of the separation $\epsilon$ and the number of photons $N$. In panel (a), the white region indicates where the MLE estimators begin to truncate negative estimates, and panel (b) displays the corresponding MSE with this truncation applied. Panel (c) presents the unbiased Cramér-Rao lower bound (CRLB). By taking the difference between the unbiased CRLB and the MSE (CRLB -- MSE), we obtain panel (d), which highlights a transition: initially, the MSE exceeds the unbiased CRLB (blue regions), but as the separation decreases and with larger bias, it drops below the CRLB (yellow regions).
IV.2 Quantum memory interaction
To characterize the mappings between the signal mode and the spin-wave mode, we can describe the mapping using Green’s functions [56], as the Raman memory equations are linear:
$B_{\mathrm{stor}}(z)=\int_{-\infty}^{+\infty}\mathrm{d}\tau\,K_{1}(z,\tau)S_{\mathrm{in}}(\tau),$ (S7)
$S_{\mathrm{out}}(\tau)=\int_{0}^{L}\mathrm{d}z\,K_{2}(z,\tau)B_{\mathrm{stor}}(z),$ (S8)
where $S_{\text{in}}(\tau)$ is the input signal mode, $B_{\text{stor}}(z)$ is the stored spin-wave mode, $S_{\text{out}}(\tau)$ is the retrieved optical mode. $K_{1}$ and $K_{2}$ are the corresponding Green’s function components that characterize the storage and retrieval processes and contain information about the corresponding control fields, such as temporal modes and intensities.
To derive the memory efficiencies, we first need to define the number of excitations $N$ in each process.
[…]
where $N_{\text{in(out)}}$ is the average number of input (output) excitations in the optical field, and $N_{\text{stor}}$ is the average number of the stored spin-wave excitations in the medium. Thus, the storage, retrieval and total memory efficiencies are given as
$\eta_{\mathrm{storage}}$ $=N_{\text{stor}}/N_{\text{in}}=1-N_{\text{tran}}/N_{\text{in}}$ (S11)
$\eta_{\mathrm{retrieval}}$ $=N_{\text{out}}/N_{\text{stor}}=N_{\text{out}}/(N_{\text{in}}-N_{\text{tran}})$
$\eta_{\mathrm{total}}$ $=N_{\text{out}}/N_{\text{in}}.$
where $N_{\text{tran}}=N_{\text{in}}-N_{\text{stor}}$ is the transmitted number of excitations. Experimentally, we used the transmitted photon counts to calculate the storage efficiency, as the number of excitations is not accessible.
The modal properties of the Raman memory can be understood by performing singular value decomposition (SVD) on the Green functions $K_{1}$ and $K_{2}$ [56]. For the storage process, $K_{1}(z,\tau)$ can be decomposed into
$K_{1}(z,\tau)=\sum_{k}\lambda_{k}\psi_{k}(z)\phi_{k}^{*}(\tau)$ (S12)
where $\lambda_{k}$ are the eigenvalues of the decomposition, sets of $\psi_{k}(z)$ and ${\phi_{k}(\tau)}$ represent the orthonormal bases of spin-waves and input temporal modes, respectively. In a single-mode quantum memory, only a specific input temporal mode $\phi_{1}(\tau)$ is efficiently coupled and stored as a corresponding spin-wave mode $\psi_{1}(z)$, while all orthogonal modes $\phi_{k}(\tau)$ for $k\neq 1$ pass through the medium unaffected. This selective interaction effectively allows the memory to function as a mode filter, storing one targeted signal while leaving others undisturbed. Crucially, the mode that is stored can be tuned by shaping the control pulse used in the Raman interaction. Adjusting the pulse’s shape, timing, and intensity modifies the Green function $K_{1}(z,\tau)$, which in turn determines the set of temporal modes ${\phi_{k}(\tau)}$ and their associated coupling strengths, represented by eigenvalues $\lambda_{k}$.
IV.3 Estimation results with different detected photon budgets
Figure S2: MLE estimation results for different total detection counts. (a), (b), and (c) correspond to total detection counts of $N=$ $2\times 10^{3}$, $10\times 10^{3}$ and $100\times 10^{3}$, respectively. The upper panels show the MLE estimators, along with the corresponding quantum bounds (blue region) and direct intensity measurement bounds (gray region). Error bars represent the estimator variances. The lower panels display the MSE (red line), with contributions from bias squared (blue region) and variance (green region).
To compare the performance of MLE estimation under different photon budgets, we present results for total detection counts of $N=$ $2\times 10^{3}$, $10\times 10^{3}$ and $100\times 10^{3}$ in Fig. S2. In all three cases, the MLE estimators closely follow the ground truth (dashed black line). The error bars fall below the intensity detection bound (gray region) but remain above the quantum limit (blue region), demonstrating improved precision over direct intensity (DI) measurements. As expected, the estimator variance decreases with increasing photon counts. The lower panels show the MSE (red line) along with its constituent components, as defined in Eq. 6 in the main text: variance (green area) and bias squared (blue area). The sum of these two contributions aligns well with the overall MSE of the estimators.
IV.4 Different retrieval modes and retrieval times
Figure S3: Performance of MLE estimation at different retrieval times. (a), (b), and (c) correspond to retrieval times of 150 ns, 200 ns, and 250 ns, respectively. In each case, the upper panels show the MLE estimates (red dots), along with the quantum bounds (blue region) and the direct intensity measurement bounds (gray region). The lower panels display the detected signal pulse sequences for two spectral lines with zero separation and zero relative phase. The input signal is shown in gray; the first red pulse indicates leaked signal, and the second red pulse corresponds to the retrieved signal. The read-in control pulse is in the HG0 mode, and the read-out control pulse is in the HG1 mode.
Figure S4: Performance of MLE estimation with different retrieval modes. (a) and (b) correspond to retrieval control modes of HG0 and HG1, respectively. The upper panels show the MLE estimates (red dots), along with the quantum bounds (blue region) and the direct intensity measurement bounds (gray region). The lower panels display the detected signal pulse sequences for two spectral lines with zero separation and zero relative phase. The input signal is shown in gray; the first red pulse indicates leaked signal, and the second red pulse corresponds to the retrieved signal. The read-in control pulse is in the HG0 mode.
Photonic quantum memories are designed to store photons over user-defined delay times. Raman quantum memories, in particular, offer additional functionality such as temporal mode conversion, where the stored signal can be retrieved into a temporal mode defined by the read-out control pulse. These combined capabilities of storage and mode conversion make Raman memories more versatile than previously demonstrated mode-filtering superresolution schemes, and could enable applications like distributed sensing networks, as discussed in the main text. Here, we demonstrate the superresolution performance of a Raman memory using different retrieval times and output temporal modes.
The maximum storage time is generally limited by thermal diffusion in the warm vapor cell: atoms carrying the spin-wave coherence drift out of the interaction region (defined by the control beam size), rendering the stored signal unrecoverable. The typical coherence time is on the order of microseconds [72]. In our experiments, we characterize the superresolution performance of the Raman memory at storage times of 150 ns, 200 ns, and 250 ns, as shown in Fig. S3. These durations are chosen based on constraints from our delay fiber length and pulse sequence design. For all three retrieval times, the superresolution performance remains comparable, as illustrated in the upper panels of the figure. The lower panels show the raw detection counts from a representative experimental run, where the input signal had zero frequency separation and zero relative phase. The retrieved signals appear as the second red pulse, occurring precisely at 150 ns, 200 ns, and 250 ns, corresponding to retrieval using an HG1 mode as the control read-out pulse.
We also performed the experiment using different retrieval modes to demonstrate the flexibility in choosing the read-out control pulse. The results are shown in Fig. S4(a) and (b), where the control read-out pulses are set to HG0 and HG1, respectively. The superresolution performance remains comparable in both cases, and the retrieved signals (the second red pulse) clearly exhibit the corresponding temporal mode shapes.
IV.5 Sources of mode crosstalk
In this work, we employed an imperfect two-mode projection method to generate our estimators, incorporating all imperfections into the parameters $\alpha$ and $\beta$ in the perturbation matrix $\bm{M}$ (Eq. 3 in the main text). Below, we discuss the primary sources of crosstalk that contribute to the mode crosstalk and the bias in the raw estimator.
IV.5.1 Memory storage efficiency
Figure S5: Effect of storage efficiency on superresolution performance. (a) Raw estimators $\hat{\epsilon}=4\sqrt{N_{1}/N_{0}}$ for various storage efficiencies, showing increased bias at higher efficiencies due to stronger temporal mode crosstalk. (b) Unbiased CRLB for the two-mode demultiplexing method at different storage efficiencies, compared to the quantum CRLB and the CRLB for the direct intensity measurement method. (c) Precision enhancement $\mathcal{F}/\mathcal{F}_{\text{DI}}$ as a function of separation $\epsilon$, demonstrating logarithmic improvement for small separations.
The primary source of temporal mode crosstalk arises from the memory’s storage interaction. Although the memory operates approximately in a single-mode regime, increasing the control field strength induces stronger AC Stark shifts. This leads to the coupling of orthogonal signal modes into the spin wave, distorting the ideal projection probabilities and introducing bias. Figure S5 illustrates how different storage efficiencies (i.e. different control read-in pulse power) affect the superresolution performance and estimation precision relative to DI methods.
Panel (a) shows the raw estimators $\hat{\epsilon}=4\sqrt{N_{1}/N_{0}}$ for different storage efficiencies. As storage efficiency increases, so does crosstalk, resulting in greater bias in the raw estimators. In panel (b), we apply maximum likelihood estimation and plot the corresponding unbiased CRLB for the two-mode demultiplexing method, alongside the quantum CRLB and the CRLB for DI methods. Higher storage efficiency leads to an increase in the CRLB, resulting in reduced estimation precision. This trend is further quantified in panel (c), where the precision improvement factor, $\mathcal{F}/\mathcal{F}_{\text{DI}}$, is plotted as a function of separation $\epsilon$. The mode demultiplexing method offers logarithmic improvements for small separations, but reaches an asymptotic limit as $\epsilon\to 0$. This limit defines the superresolution parameter [43], which is plotted against storage efficiency in panel (d). As storage efficiency decreases, the superresolution parameter increases logarithmically. However, very high precision improvement using very low storage efficiencies is impractical in experiments, as fewer signal photons are retrieved, constant background and control field leakage reduce the signal-to-noise ratio, potentially increasing the observed crosstalk and estimation bias.
IV.5.2 Control leakage
Another major limiting factor contributing to excess crosstalk in our experiment is control field leakage, which introduces a constant background noise. Due to the small frequency separation (9.2 GHz) between the signal and control fields, completely filtering out the control field is technically challenging. As a result, residual control photons are detected and contribute equally to both HG0 and HG1 projection measurement counts, adding uniform noise across all separations that distorts the projection probabilities.
To quantify this effect, we simulate how constant background counts influence the bias of the raw estimators and the precision improvement of MLE estimators. In this simulation, we fix the storage efficiency at 40% and introduce a constant leakage level ranging from 0 to 2% of the total detected counts. The results are shown in Fig. S6. Panel (a) shows that increasing control leakage leads to larger bias in the raw estimator, similar to the effect of increasing storage efficiency. The rate of bias increase slows as the leakage grows uniformly. Panel (b) presents the corresponding precision improvement, which diminishes with higher background noise due to the increasing bias.
Figure S6: Effect of control leak counts on estimation bias and precision. (a) The bias in the raw estimators increases with control leakage, expressed as a percentage of the total detection counts. (b) Precision enhancement of the mode filtering method decreases as background noise increases. Storage efficiency is fixed at 40%.
IV.5.3 Others
Additional sources of mode crosstalk include detector dark counts, signal background from incomplete extinction of the electro-optic modulator, and four-wave mixing (FWM) noise generated during the Raman interaction. In our experiment, we minimized the signal background using a Pockels cell and characterized both the dark counts and residual background, which were subtracted from the measured data. Furthermore, we applied appropriate detunings to the Raman transition to naturally suppress FWM noise, as described in Ref. [55]. Thus, the contribution of these sources is minimal.
IV.6 Precise pulse carving
To improve our pulse, carving system, we applied frequency response correction to our electronic system. In a linear time-invariant (LTI) system, where our electronic system can be approximated as such, the electric signal output $h(t)$ of the system to an arbitrary electric signal input function $x(t)$ can be approximately described by a linear response function $R(t-t^{\prime})$ as
[…]
In the frequency domain, the frequency response function $\tilde{R}(\omega)$ can be calculated as,
$\tilde{R}(\omega)=\frac{\tilde{h}(\omega)}{\tilde{x}(\omega)}.$ (S14)
Once the response function is known, a target signal $h^{\prime}(t)$ can be precisely generated by using the response-corrected input signal $x^{\prime}(t)$:
$x^{\prime}(t)=\mathcal{F}^{-1}\left[\frac{\mathcal{F}(h^{\prime}(t))}{\tilde{R}(\omega)}\right].$ (S15)
where $\mathcal{F}$ denotes the Fourier transform and $\mathcal{F}^{-1}$ denotes the inverse Fourier transform.
To demonstrate our high-fidelity optical pulse carving system, we present the averaged and normalized intensity profiles of the HG0 and HG1 optical control pulses, as measured by a photodiode and an oscilloscope (1024 averages). Fig. S7(a) and (b) show the experimentally measured pulse shapes (red) compared against the theoretical intensity distributions (black dashed lines). The discrepancies between the theoretical and experimental data are plotted in the lower panels. The deviation between the measured and theoretical profiles is small, typically not exceeding 2% relative to the maximum.
Figure S7: Intensity profiles of experimental HG control pulses. (a) and (b) show the measured intensity profiles of the HG0 and HG1 control pulses, respectively. In the upper panels, the experimentally measured normalized profiles (solid red) are compared with the theoretical normalized HG profiles (dashed black). The lower panels display the discrepancies between the experimental and theoretical profiles.