### v2

# Super-resolving frequency measurement with mode-selective quantum memory

Abstract
High-precision optical frequency measurement is indispensable to modern science and technology, yet conventional spectroscopic techniques struggle to resolve sub-linewidth spectral features. We introduce a unique platform for super-resolving frequency estimation utilizing a mode-selective atomic Raman quantum memory implemented in warm cesium vapor. By precisely engineering the light-matter interaction, our memory coherently stores the optimal temporal mode with high fidelity and retrieves it on-demand, realizing a mode crosstalk as low as 0.34%. To estimate the separation between two spectral lines, we experimentally measure the mean squared error of the frequency estimate, achieving a sensitivity of 1/20 of the linewidth with a (34±4)-fold enhancement in precision over direct intensity measurements. This enhanced frequency resolution, combined with the memory’s on-demand storage, retrieval, and mode conversion capabilities, paves the way for multifunctional memory-based time-frequency sensors and their integration within quantum networks.
I Introduction
The time-frequency (TF) degree of freedom of light underpins a broad range of applications spanning high-resolution spectroscopy [1], precision timekeeping [2, 3], ultrafast optics [4] and emerging quantum technologies [5, 6, 7, 8]. In spectroscopy, for example, resolving spectral lines is essential for probing the atomic and molecular properties of matter. Frequency measurements are also crucial for quantum metrology and sensing [9, 10, 11], as they largely rely on measuring the transition between quantized energy levels. However, the precision of these optical measurements has long been constrained by instrumental limitations. Every spectral feature has an intrinsic linewidth set by the Fourier limit - a lower limit being the inverse of the measurement time or the signal’s temporal duration. The resolving power of conventional spectrometers is compounded by the Rayleigh criterion [12], which defines the minimal resolvable separation between two spectral lines. As the separation shrinks, the uncertainty in their resolution escalates rapidly. This phenomenon, often termed Rayleigh’s curse, poses a critical barrier, particularly when probing weak or photon‐limited signals.
The ultimate precision in resolving spectral features is bounded by the quantum mechanical nature of the optical field [13]. From quantum estimation theory, Tsang, Nair, and Lu have shown that Rayleigh’s criterion is not a fundamental limit, but rather an artifact of the direct intensity measurement (DI) strategy - with coherent measurements an arbitrarily small separation can be resolved with a constant, fine precision [14]. To circumvent the classical resolution limit, implementing a coherent mode filter to select the optimal mode basis before detection, instead of DI measurements, is essential [15, 16, 17, 18, 19, 20]. The practical realization of this advantage, however, relies critically on the accuracy of mode filtering, as experimental imperfections like mode crosstalk and detector noise substantially diminish the achievable precision [21, 22, 23, 24, 25, 26]. The high-fidelity coherent mode filtering is generally needed for broader quantum information processing and quantum metrology technologies [6, 7, 27, 28, 29]. Looking ahead, the future quantum networks demand sensor nodes with such capability of coherent TF processing, ideally integrated with on-demand buffering for signal synchronization [30, 31], coherent bandwidth and frequency conversion to interface different physical platforms and network channels [32, 33, 34] - all while preserving quantum coherence. These integrated functionalities are essential for building robust, distributed quantum-enhanced sensors [35] capable of dynamically adjusting to varying environmental conditions across a network [36, 37, 38, 39].
In TF super-resolution, quantum pulse gates (QPGs), which use nonlinear waveguides for mode-selective frequency conversion, have been deployed to resolve temporal and spectral separations for ultrafast pulses with hundreds of GHz bandwidth [40, 41]. While QPGs demonstrate programmable mode selectivity [42], they inherently lack on-demand storage and buffering capabilities. In the narrow-band frequency regime, time-inversion interferometry using gradient echo memory has demonstrated sub-Rayleigh resolution for tens of kHz bandwidth pulses [43], but they are limited in the ability to select arbitrary modes and require cumbersome magnetic field gradients and cryogenics. Additionally, there have also been works focusing on frequency super-resolution across tens of GHz bandwidth using dispersion engineering and electro-optic modulators [44, 45]. However, no existing platform delivers high‐precision super‐resolution in the MHz to GHz bandwidth together with on‐demand storage, retrieval and user-defined mode selectivity.
In this work, we introduce a unique scheme of TF super-resolution via high-fidelity coherent mode filtering in atomic Raman quantum memory. Photonic quantum memories have been widely studied for absorbing and re-emitting photonic states on demand, with applications in quantum networks, communication and computing [46, 47, 48, 49, 50, 51, 52]. Implemented in a warm cesium vapor, our platform utilizes a three-level atomic system where a strong control field with tailored temporal profile coherently maps an incoming signal field onto a collective atomic spin-wave coherence via stimulated Raman absorption, achieving high-fidelity mode-selectivity up to 99.6% for orthogonal Hermite-Gaussian (HG) temporal modes. Leveraging this user-defined coherent mode filter, we store the optimal signal mode containing the information of spectral line separation, retrieve it on demand, and apply maximum-likelihood estimation to the retrieved photon statistics to extract frequency separations. Focusing on sub-linewidth frequency separation under various detected photon budgets (from $2\times 10^{3}$ to $1\times 10^{5}$), our platform consistently outperforms DI methods, achieving high precision enhancements. Operating in the MHz to GHz bandwidth, our memory-based platform extends the toolbox of TF metrology and provides integrated functionalities encompassing mode filtering, buffering, and shaping. These capabilities are ideally suited for next-generation quantum sensor nodes and their deployment in quantum networks.
II Results
II.1 Super-resolving measurement in frequency domain
We frame the task of resolving two closely spaced spectral features as an estimation of their normalized frequency separation, $\epsilon=\Delta\omega/\sigma$. Here, $\Delta\omega$ denotes the frequency difference between the two spectral lines centered at $\omega_{0}\pm\Delta\omega/2$, and $\sigma$ is the spectral linewidth. We model the sources as two mutually incoherent emitters of equal intensity, each described by a Gaussian spectral amplitude $\psi(\omega)=(2\pi\sigma^{2})^{-1/4}\exp(-\omega^{2}/4\sigma^{2})$. Conventional spectroscopy measures the power spectrum,
$S(\omega|\epsilon)=\frac{1}{2}\left[\left|\psi\left(\omega-\omega_{0}-\frac{\epsilon\sigma}{2}\right)\right|^{2}+\left|\psi\left(\omega-\omega_{0}+\frac{\epsilon\sigma}{2}\right)\right|^{2}\right].$
The separation $\epsilon$ is then estimated by fitting this model to the measured spectrum. The amount of information about $\epsilon$ extractable from the measurement outcomes is quantified by the Fisher information (FI) per detected photon
$\mathcal{F}_{\text{DI}}(\epsilon)=\int_{-\infty}^{\infty}d\omega\frac{1}{S\left(\omega|\epsilon\right)}\left(\frac{\partial S\left(\omega|\epsilon\right)}{\partial\epsilon}\right)^{2},$
which vanishes in the limit $\epsilon\to 0$ [14]. The precision, quantified by the variance of an unbiased estimator $\hat{\epsilon}$, is lower-bounded by the Cramér–Rao lower bound (CRLB) $\text{Var}(\hat{\epsilon})\geq 1/[N\mathcal{F}(\epsilon)]$, where $N$ represents the number of detected photons. As $\epsilon\to 0$, the two Gaussian lines increasingly overlap and the estimator variance diverges to infinity, making the estimation of arbitrarily small separations infeasible.
An optimal measurement, in contrast, projects the signal onto an orthonormal Hermite–Gaussian (HG) mode basis with coherent filtering. Experimentally, projecting the input signal onto the HG basis in the frequency domain is equivalent to filtering its temporal profile - the Fourier transform of its spectrum - with the corresponding temporal HG basis. This mode basis exhibits a constant FI of $\mathcal{F}_{\text{HG}}(\epsilon)\approx 1/4$ that saturates the quantum FI - the ultimate limit of precision attainable over all possible measurements for a given input quantum state [14]. As the HG measurement is particularly advantageous for small separations, we will focus on the precision and ability to resolve small separations under limited photon budgets in realistic experiments. For small separations, the projection onto the $\text{HG}_{0}$ and $\text{HG}_{1}$ mode contributes most of the FI. The ideal projection probabilities onto HG0 and HG1 modes defines a raw estimator $\hat{\epsilon}_{\text{raw}}=4\sqrt{N_{1}/N_{0}}$, where $N_{0}$ and $N_{1}$ are the experimentally measured counts of HG0 and HG1 projections.
In practical implementations, mode crosstalk modifies the ideal projection probabilities. We can model this effect between the first two HG modes by describing how the ideal probability vector, $\bm{p}$, is perturbed into the measured probability vector, $\bm{\tilde{p}}$. This transformation is defined by a coupling matrix $\bm{M}$, as
$\bm{\tilde{p}}=\bm{Mp}=\begin{pmatrix}\alpha&1-\beta\\ 1-\alpha&\beta\end{pmatrix}\begin{pmatrix}P(0|\epsilon)\\ P(1|\epsilon)\end{pmatrix}$, (1)
where $\alpha,\beta\in[0,1]$ quantify the mode filtering fidelity. The resulting perturbed probability $\bm{\tilde{p}}=\begin{pmatrix}\tilde{P}(0|\epsilon)\\ \tilde{P}(1|\epsilon)\end{pmatrix}$ is then normalized. Assuming low mode crosstalk $\alpha\approx 1$ and $\beta\approx 1$, we derive the FI for small separations as (see Supplementary Information for details)
$\mathcal{F}(\epsilon)\approx\frac{1}{4\left[(1-\alpha)/(\epsilon/4)^{2}+1\right]}.$ (2)
The FI in the presence of mode crosstalk is highly sensitive to the leakage from the HG0 mode into the HG1 mode, $1-\alpha$. If $(\epsilon/4)^{2}\ll 1-\alpha$, the FI decreases significantly from the ideal case, highlighting the importance of low-crosstalk mode filtering in resolving very small separations.
To estimate $\epsilon$ in the presence of crosstalk, we employ maximum likelihood estimation (MLE) using the perturbed projection probabilities onto the first two HG modes, $\tilde{P}(0|\epsilon)$ and $\tilde{P}(1|\epsilon)$. The MLE finds the parameter value that maximizes the likelihood function $\mathcal{L}(\bm{x}|\epsilon)=\prod_{i=1}^{N}\tilde{P}(x_{i}|\epsilon)$ for observing the measured data $\bm{x}=\{x_{1},x_{2},\ldots,x_{N}\}$. For small separations, the MLE estimator can be written as
$\hat{\epsilon}_{\text{MLE}}=\underset{\epsilon}{\mathrm{argmax}}[N_{0}\ln{\tilde{P}(0|\epsilon)}+N_{1}\ln{\tilde{P}(1|\epsilon)}],$ (3)
subject to the parameter space constraint $\Theta=\{\epsilon\in\mathbb{R}\mid\epsilon\geq 0\}$. The MLE is asymptotically unbiased and efficient in the limit $N\to\infty$, implying its variance approaches the CRLB.
In real experiments with finite statistics, the MLE exhibits a non-zero bias, particularly at small separations. This bias arises from both higher-order asymptotic terms [53] and the non-negativity parameter space constraint $\Theta$. Due to shot noise, the experimentally observed normalized counts $\bm{f}=\begin{pmatrix}N_{0}/N\\ N_{1}/N\end{pmatrix}$ may lie outside the physically valid range. When this occurs, the likelihood function $\mathcal{L}(\bm{x}|\epsilon)$ takes its maximum at the boundary of the parameter space constraint (i.e. $\epsilon=0$), leading to a positively skewed distribution of the estimator. This skewing effect results in an unavoidable bias for $\epsilon\to 0$, with the magnitude of the bias determined by the estimator’s distribution. The bias diminishes with higher photon counts $N$, as statistical fluctuations on $N_{0}$ and $N_{1}$ decrease, making the estimator’s distribution at the boundary less probable (see Supplementary Information for simulation results).
We assess our measurement performance using the mean squared error (MSE), $\text{MSE}(\epsilon,N)=\langle(\hat{\epsilon}-\epsilon)^{2}\rangle$, incorporating both the variance and the squared bias of the estimator, $\text{MSE}(\epsilon,N)=\text{Var}(\epsilon)+b(\epsilon,N)^{2}$. The MSE is lower bounded by the CRLB with a bias term [54]
$\text{MSE}(\epsilon,N)\geq\frac{[1+b^{\prime}(\epsilon,N)]^{2}}{N\mathcal{F}(\epsilon)}+b(\epsilon,N)^{2},$ (4)
where $b^{\prime}(\epsilon,N)$ represents the derivative of $b(\epsilon)$ with respect to $\epsilon$. Only as $b(\epsilon)\rightarrow 0$ does the MSE converge to the standard unbiased CRLB, $\text{Var}(\hat{\epsilon})\geq 1/[N\mathcal{F}(\epsilon)]$. From the MSE, we quantify the sensitivity of the apparatus as the minimal resolvable separation $\epsilon_{\text{min}}$ for which the parameter-to-error ratio (PER) of the estimate,
$\text{PER}(\epsilon,N)=\epsilon^{2}/\text{MSE}(\epsilon,N),$ (5)
is greater than 1 [22]. This minimal resolvable separation defines a modified Rayleigh criterion for resolving spectral lines under certain detected photon budgets. In this work, MSE and PER serve as the figures of merit to quantify the performance of our experimental platform.
II.2 Mode-selective Raman quantum memory
[…]
In the low-coupling regime (i.e., low storage efficiencies), the memory operates in a single-mode fashion: the stored temporal mode is directly determined by the temporal profile of the control pulse [55] (see Supplementary Information for details). Fig. 1(B) illustrates two representative cases: when both signal and control are in the $\mathrm{HG}_{0}$ mode (label 1), the signal is efficiently stored and retrieved; conversely, when the signal is in $\mathrm{HG}_{0}$ but the control is in $\mathrm{HG}_{1}$ (label 2), storage and retrieval are strongly suppressed. Experimental results for these scenarios are shown in Fig. 1(C). To quantify the mode selectivity, we measured the total efficiencies for all combinations of signal and control pulses prepared in HG0 and HG1 modes. Fig. 1(D) shows the measured mode crosstalk matrix with high mode selectivity: $(99.61\pm 0.48)\%$ when storing $\mathrm{HG}_{0}$ signals with $\mathrm{HG}_{0}$ control, and $(0.39\pm 0.02)\%$ with $\mathrm{HG}_{1}$ control. These results were obtained using read-in and read-out control pulses with energies of $125\pm 1$ pJ, yielding total efficiencies of approximately 11% along the diagonal terms. At higher coupling strengths (e.g., with stronger control pulses), this single-mode behavior degrades due to the increased storage of temporal modes orthogonal to that of the intended control field. This highlights an inherent trade-off between mode selectivity and overall memory efficiency. By achieving low crosstalk and high mode selectivity in the low-coupling regime, our Raman memory provides a robust and effective platform for the super-resolution task.
II.3 Experimental superresolution with mode-selective quantum memory
Figure 2: Experimental estimation results. This graph compares the maximum likelihood estimators (red markers) with raw estimators (yellow markers), both derived from $10\times 10^{3}$ total detected photons. The dashed black line indicates the ground truth separation. The shaded regions illustrate the theoretically predicted standard deviations from the quantum CRLB (blue) and the direct intensity measurement CRLB (gray).
To perform estimation experiments using the mode-selective Raman quantum memory described above, we first prepared the signal as an incoherent mixture of two Gaussian spectral lines, each with a linewidth of $\sigma=5.30$ MHz, and varied their normalized separation over 21 values from 0 to $1$ in steps of $0.05$ (see Methods). For each separation, we performed two measurements by storing the signal with either an $\mathrm{HG}_{0}$ or an $\mathrm{HG}_{1}$ control read-in pulse. In both cases, the stored signals were then retrieved after 200 ns using an $\mathrm{HG}_{1}$ pulse and detected with SNSPDs. At $\epsilon=0$, the storage efficiency and total internal efficiency for the $\text{HG}_{0}$ control read-in were measured to be $(39.5\pm 0.6)\%$ and $(14.6\pm 0.1)\%$ respectively, using control pulses with energies of $130\pm 1$ pJ. For each separation, we recorded approximately $N\approx 2\times 10^{5}$ detector clicks in total.
To estimate the frequency separation $\epsilon$, we first calibrated the experimental system with $1.6\times 10^{5}$ detected counts per separation. By fitting the experimental model to this calibration data, we determined the measured crosstalk from $\text{HG}_{0}$ to $\text{HG}_{1}$ in the estimation experiment to be 0.34%. It is noteworthy that this is a separate experiment from the mode filtering demonstration in the previous section, accounting for the slight difference in the reported crosstalk. With the calibrated model, in each run we performed MLE to compute the estimate $\hat{\epsilon}_{\text{MLE}}$ based on Eq. 3 with registered counts $(N_{0},N_{1})$. We then obtained the MSE for each separation by averaging $(\hat{\epsilon}_{\text{MLE}}-\epsilon)^{2}$ over 50 bootstrapped experiments with the same $N$ drawn from the full click dataset. The uncertainty in the MSE was evaluated by repeating this procedure 10 times.
Figure 2 shows the estimation results with estimator standard deviations for $N=10\times 10^{3}$ detected counts. The complete estimation results for various detection counts, detailing the contributions from estimator bias and variance, are presented in the Supplementary Information. While the raw estimators $4\sqrt{N_{1}/N_{0}}$ have low variance, they suffer from a systematic bias caused by mode crosstalk and control field leakage. In contrast, the MLE exhibits substantially reduced bias. For the smallest separations ($\epsilon=0-0.2$), the MLE estimators display lower bias but higher variance, reflecting the trade-off between estimator variance (precision) and bias (accuracy) described in Eq. 4. Crucially, all MLE error bars are smaller than the DI bound, demonstrating enhanced precision over DI detection. To validate the platform’s unique functionality of on-demand storage, retrieval and coherent mode conversion, we also performed super-resolving estimation with storage times from $150$ ns to $250$ ns and with $\mathrm{HG}_{0}$ and $\mathrm{HG}_{1}$ retrieval modes. The results are included in the Supplementary Information.
The residual MLE bias near $\epsilon=0$ is mainly attributed to the non-negativity constraint on $\epsilon$. This estimator bias is unavoidable under limited photon budgets and depends on the number of detected photons. Fig. 3(A) plots $\mathrm{MSE}\times N$ together with the corresponding CRLBs for comparison across different detected counts. For $\epsilon<0.5$, the measured MSE is generally smaller than the DI CRLB, indicating improved precision. However, at low photon numbers (e.g. $N=2\times 10^{3}$), the MSE falls below the mode-filtering CRLB for $\epsilon<0.15$ due to estimation bias. As the photon count increases to $100\times 10^{3}$, the MSE approaches the CRLB as the bias decreases, which is consistent with our earlier discussion. Thus, for limited photon budgets the MLE incurs a small bias at small separations, but its contribution to the MSE becomes negligible as $N$ increases, allowing reliable, high-precision estimation at larger photon numbers.
Figure 3: Estimation performance under different photon budgets. (A) MSE scaled by photon number NN. Experimental results for $N=2\times 10^{3}$ (orange triangles), $10\times 10^{3}$ (green squares), and $100\times 10^{3}$ (blue circles) approach the theoretical CRLB of our setup (solid red line) as NN increases. The quantum CRLB (blue dash-dotted line) and the direct intensity detection CRLB (gray dashed line) are shown for reference. Error bars are from 10 bootstrapped results. (B) Parameter-to-error ratio for different detection counts. Experimental data (markers) are plotted alongside their corresponding theoretical values (solid lines). The dashed line at 0 dB marks the sensitivity threshold.
II.4 Sensitivity and precision enhancement benchmarking
To benchmark the sensitivity of our measurement scheme, we examine the minimum resolvable separation based on the PER ratio, as defined in Eq. 5. Fig. 3(B) shows the PER as a function of separation for different total detection counts. With $100\times 10^{3}$ detected photons, we achieve $\mathrm{PER}=4.4\pm 0.5$ dB at $\epsilon=0.05$, corresponding to resolving spectral lines separated by just $265$ kHz for a linewidth of $5.30$ MHz. In contrast, for lower photon numbers ($N=2\times 10^{3}$ and $10\times 10^{3}$), $\mathrm{PER}<0$ dB at $\epsilon=0.05$, indicating that such small separations cannot be distinguished from zero separation at these photon levels. The theoretical PER values (solid lines in Fig. 3(B)) show good agreement with the experimental results.
In addition to sensitivity, we evaluate the precision enhancement enabled by our memory-based mode filtering scheme against DI methods. We quantify the enhancement via the superresolution parameter $\mathfrak{s}$ introduced by Mazelanik et al. [43], which is defined as the ratio of the FI for the two methods in the limit of vanishing separation $\epsilon\to 0$
$\mathfrak{s}=\lim\limits_{\epsilon\rightarrow 0}(\mathcal{F}/\mathcal{F}_{\text{DI}}).$ (6)
This parameter defines the precision enhancement achievable with an infinite photon budget, as a truly unbiased estimator is only realizable when $N\to\infty$. This parameter can be used to assess and compare the performance of various super-resolving measurement schemes, with its value primarily determined by mode crosstalk. Figure 4(A) shows how the theoretical enhancement $\mathcal{F}/\mathcal{F}_{\mathrm{DI}}$ (red curve) varies with $\epsilon$. In the limit $\epsilon\rightarrow 0$, the ratio asymptotically approaches $\sim 37$ for our platform. This value sets a new benchmark for frequency-separation estimation, surpassing previously reported results [40, 43, 45], as shown in the comparative graph in Fig. 4(B). For comparison, we also show the performance of DI methods — including quantum-memory temporal imaging [43], a Fourier transform spectrometer and grating spectrometer - all of which exhibit $\mathfrak{s}<1$. Among techniques spanning various bandwidth regimes, our Raman memory uniquely operates in the MHz to GHz range, set by the hyperfine splitting and the memory storage time.
It should be noted that under finite statistics, the experimental precision enhancement may deviate from the theoretical values due to estimator bias. In practice, when the estimator exhibits negligible bias, the enhancement can be approximated using the ratio of the measured mode filtering MSE to the CRLB of DI methods. Shown as data points in Fig. 4(A), these ratios align well with the trend of the theoretical prediction, subject to experimental imperfections and fluctuations. However, at very small separations, the estimator bias can cause the measured MSE to fall below the CRLB, resulting in calculated ratios that exceed the theoretical values. In this region, we expect both the mode filtering method and the DI detection method exhibit estimator biases, making the determination of realistic precision enhancement more complex [21]; a detailed analysis will be explored in future work. Our experiment demonstrates a practical $(34\pm 4)$-fold improvement at $\epsilon=0.05$ and a $(28\pm 6)$-fold improvement at $\epsilon=0.1$ using $100\times 10^{3}$ detected photons, confirming the feasibility and effectiveness of our approach. Furthermore, we expect DI measurements, which inherently have larger estimator variances from the FI analysis, require higher number of photons to achieve nearly unbiased estimation, whereas the optimal mode filtering maintains lower variance even in photon-limited applications.
Figure 4: Precision enhancement benchmarking. (A) Precision enhancement as a function of spectral line separation $\epsilon$. The theoretical precision enhancement (red curve) is plotted with all experimental results for various photon counts: $N=2\times 10^{3}$ (orange triangles), $10\times 10^{3}$ (green squares), and $100\times 10^{3}$ (blue circles). (B) Comparison of superesolution parameters for various schemes. This panel benchmarks the performance of our system against other super-resolution techniques and DI methods. QPG: quantum pulse gate [40]; PuDTAI: pulsedivision time-axis-inversion [43]; SUSI: super-resolution via spectral inversion [45]; QMTI: quantum-memory temporal imaging [43]; FT: Fourier transform spectrometer (Bruker IFS 125HR); Grating: grating spectrometer of grating length 1 cm and grating period of 1200 mm.
III Discussion
We have demonstrated a clear improvement in the precision of spectral line separation estimation. Our system achieves a practical sensitivity sufficient to resolve separations as small as 1/20 of the individual signal linewidth, providing a $(34\pm 4)$-fold precision enhancement over direct-intensity methods at this separation. Among TF superresolution techniques [40, 43, 45], our Raman memory operates over a unique bandwidth, from MHz to GHz, where no classical strategy is available. Going beyond resolving two spectral lines, our platform’s fully programmable temporal mode filtering also opens the door to resolving more complex spectral features. The user-defined mode filtering is pivotal for multi-parameter estimation, where the optimal measurement basis can be complicated depending on the task [56, 57, 58, 59]. Such optimal measurement bases are ideally realized through TF mode sorting onto multiple temporal modes [42], which could be implemented using multiple cascaded memories or a loop architecture with a single memory.
As indicated in Eq. 2, achieving even higher precision hinges on reducing mode crosstalk, which primarily stems from the memory storage process. Crucially, there is a trade-off between mode crosstalk and storage efficiency, which arises from the low-coupling requirement for the memory’s single-mode behavior (see Supplementary Information for simulation results). To address this limitation, our recently developed efficiency enhancement via light-matter interference (EEVI) protocol [60] offers a path to realizing high-efficiency storage and retrieval while preserving single-modeness. Cavity-enhanced Raman memories provide a similar solution [61, 62], though their bandwidth is constrained by cavity finesse. Alternatively, optimal control techniques can tailor pulse shapes to maximize storage efficiency for the target mode while suppressing orthogonal modes, thereby minimizing crosstalk [63, 64]. These strategies offer promising pathways toward overcoming the trade-off between storage efficiency and crosstalk.
Despite the importance of precision enhancement, the practical advantage of the super-resolving measurement, quantified by FI per input photon, is greatly impacted by the photon loss of the setup. To date, all experimental platforms for TF super-resolution suffer from low end-to-end efficiency. In our current setup, we measure an overall efficiency of approximately 0.3%, despite achieving an internal total efficiency of $(14.6\pm 0.1)\%$. This disparity is caused by the stringent filtering needed to suppress the co-propagating control field, detuned by just 9.2 GHz. Achieving a high extinction ratio of $\sim 4\times 10^{7}$ requires a series of lossy optics, including a calcite beam displacer, a Glan-Taylor polarizer, and three double-passed Fabry–Pérot etalons. In contrast to Raman memories, ladder-type quantum memories, such as fast-ladder memory [65] or off-resonant cascaded absorption memory (ORCA) [66, 67], exploit a higher-lying excited state and widely separated, counter-propagating signal and control fields. This configuration, allowing efficient suppression of control leakage without complex filtering, has demonstrated noise per signal photon as low as $10^{-5}$ [66] with end-to-end efficiencies up to 35% [65]. Integrating our high-fidelity mode filtering with these architectures holds the potential for realizing unconditional TF super-resolution.
In conclusion, we have demonstrated a frequency superresolution protocol using a mode-selective Raman quantum memory. The technology developed here holds promise for diverse applications, including high-precision clock synchronization [68], spacetime positioning [69, 70], photon dose-limited metrological tasks [7], and processing of TF encoded quantum information [6, 28, 27, 29]. Future works will focus on further improving precision through the implementation of the EEVI protocol and control optimization techniques. We will also explore the potential of ORCA memories to achieve unconditional super-resolving measurements. Furthermore, we hope to demonstrate the platform’s capabilities for multi-parameter estimation and its integration into quantum sensor networks. Ultimately, the unique combination of high-fidelity mode filtering, on-demand storage and mode conversion capabilities positions our platform as a multifunctional tool for advanced TF metrology, paving the way for integrated quantum sensors tailored to photon-limited applications and distributed sensing networks.
IV Methods
Experimental setup
Figure 5: Schematic of experimental setup. (A) Experimental setup. The apparatus consists of three main sections: pulse generation, memory interaction, and filtering and detection. Both signal and control pulses originate from a continuous wave (CW) laser at the control-field frequency. An electro-optic modulator (EOM), EOM 1, first carves the CW light into pulses. A 90:10 beam splitter (BS) then splits the laser into a high-power control path (90%) and a low-power signal path (10%). The laser in the signal path is later shifted to the signal frequency via EOM 2. The two beams are recombined via a beam displacer (BD) before entering the memory cell for storage and retrieval. Post-retrieval filtering comprises a Glan-Taylor polarizer (GTP), three double-passed etalons, and a band-pass filter (BP). The signal is finally routed to fiber beam-splitters and four superconducting nanowire single-photon detectors (SNSPDs) for detection. Employing multiple detectors helps mitigate detector dead time effects and enables higher overall count rates. (AWG: arbitrary waveform generator; PC: Pockels cell; PBS: polarizing beamsplitter; M: mirror; DM: dichroic mirror; HWP: half-wave plate; WPs: a half-wave plate and a quarter-wave plate.) (B) Pulse sequences. The diagram illustrates the timing of the signal and control pulses, along with the gating windows of the optical pumping and the Pockels cell.
Our Raman memory setup is shown in Fig. 5(A). The storage medium consists of a 75 mm-long cesium (Cs-133) vapor cell with a three-layer $\mu$-metal magnetic shielding. We heated the cell to $105\pm 1$ °C using a quad-twisted cryogenic wire, achieving an effective optical depth of $(4.8\pm 0.1)\times 10^{3}$. The atoms were initialized to the initial state $|g\rangle$ by an external cavity diode laser (ECDL, Toptica DL100) driving the D1 transition ($6^{2}\textbf{S}_{1/2}\rightarrow 6^{2}\textbf{P}_{1/2}$). An electro-optical modulator (EOM 3, EOSpace 850UL) gated the pump laser, switching it off during the experimental window. We achieved a pumping efficiency of $(99.6\pm 0.1)$%.
We operated our memory with a detuning $\Delta=2\Delta_{\text{hf}}$ to suppress four-wave-mixing noise [71]. Both the signal and control fields were derived from a single continuous wave (CW) ECDL laser (Toptica DL Pro) operating at the control frequency (351.7122 THz). An arbitrary waveform generator (AWG, Tektronix AWG70001A) drove the electro-optical modulator EOM 1 (Sacher Lasertechnik AM830PF) to carve the CW laser into pulses. The laser was then split using a 90:10 beam splitter, directing 90% of the optical power to the control field path. The remaining 10% went to EOM 2 (New Focus 4851), which generated a sideband at the signal frequency (351.7030 THz). We selected this sideband as the signal field by using a Fabry-Pérot etalon with a free spectral range (FSR) of 36.8 GHz. The signal field was delayed by 437 ns using a long fiber to achieve temporal synchronization with the control pulses, ensuring their temporal overlap for memory storage and retrieval. Before entering the memory, the signal field also passed through a Pockels cell to remove residual CW background in the retrieval window. The polarization of the signal and control fields was set to orthogonal for memory interaction, allowing their combination using a beam displacer before being focused into the vapor cell.
[…]
After the memory, we suppressed the control and pump fields using a Glan-Taylor polarizer, followed by three double-passed etalons (two with FSR of 18.4 GHz and one with FSR 103 GHz) and a bandpass filter (central wavelength: 850 nm, FWHM Bandwidth: 10 nm). Finally, we split the signal into four equal parts and detected them using four superconducting nanowire single-photon detectors (SNSPDs, Photon Spot). Employing multiple detectors helps mitigate detector dead time effects and enables higher overall count rates. The detected counts were registered using a time tagger (Swabian TimeTagger20).
Signal preparation
The signal in our separation estimation task is an incoherent mixture of two Gaussian spectral lines. Its frequency-domain representation can be written as
$\Psi(\omega)=\frac{1}{\sqrt{2}}$ $\left[\exp\left({-i\frac{\phi}{2}}\right)\right.\psi(\omega-\omega_{0}-\frac{\epsilon\sigma}{2})$ (7)
$+\left.\exp\left({i\frac{\phi}{2}}\right)\psi(\omega-\omega_{0}+\frac{\epsilon\sigma}{2})\right],$
where $\phi\in(-\pi,\pi]$ is a random phase between the two spectral lines. Here, we carved the temporal profile of our signal field as the inverse Fourier transform of Eq. 7,
$\Psi(t)=A\cos\left(\frac{\epsilon\sigma t-\phi}{2}\right)\exp(-t^{2}\sigma^{2}),$ (8)
where $A$ is the amplitude. The signal pulse has a spectral width of $\sigma=33.3$ Mrad/s (5.30 MHz). In our experiment, we generated signals with separation parameters $\epsilon$ from 0 to 1 with an increment of 0.05 (0.265 MHz). To introduce the incoherence of the two spectral lines, we prepared the signal in a mixture of four different relative phases, $\phi=-\pi/2,0,\pi/2,\pi$, for each separation.
One key requirement for this experiment is the accurate generation of signal and control pulses. To ensure high-fidelity pulse carving, we characterized the frequency response of our electronic pulse generation system, which comprises the AWG, an RF splitter, and two RF amplifiers that drive the EOM 1. By measuring the system’s frequency response function, we applied a correction to the input signals of the AWG, compensating for frequency-dependent variations in the RF components. This correction ensures uniform amplification across all frequency components, minimizing distortions in the carved pulses. The measured intensity pulse shapes of HG0 and HG1 are presented in the Supplementary Information.
Data collection and analysis
[…]
To obtain the final detected counts, we first identified a reference pulse count that yielded approximately $N$ = $2\times 10^{3}$, $10\times 10^{3}$, or $100\times 10^{3}$ total retrieval counts (with a variation less than 5%) of the signals across all combinations of the two control read-in pulses and four phases. For each separation, we used the reference counts to randomly sample a subset that yielded $N$ retrieved counts from the data files. The corresponding noise counts were subtracted from the retrieved counts to correct the background contributions. This random sampling approach, combined with the use of reference counts, was employed to mitigate the impact of signal power fluctuations on the final estimation process. The retrieval counts acquired for HG0 and HG1 storage, $N_{0}$ and $N_{1}$ respectively, were used directly to compute the raw estimator. We repeated the entire sampling and analysis procedure 50 times using bootstrapping to estimate the variances and MSEs of the estimators. To obtain the error bars for the MSE shown in Fig. 3, we repeated the bootstrapping process 10 times to get 10 datasets and calculated the standard deviations of the MSE.
System calibration
To characterize the crosstalk matrix $\bm{M}$ of the experimental system and the perturbed HG projection probabilities $\tilde{P}(0|\epsilon)$ and $\tilde{P}(1|\epsilon)$ in Eq. 3, we used a dataset for separations $\epsilon$ from 0 to 1 with $N=1.6\times 10^{5}$ detected photon counts for each separation. For each separation, we normalized the retrieved counts $N_{0}$ and $N_{1}$ corresponding to the two HG projections to obtain the relative frequencies $f_{0}=N_{0}/(N_{0}+N_{1})$ and $f_{1}=N_{1}/(N_{0}+N_{1})$. The values of $\alpha$ and $\beta$ were estimated by minimizing the least square cost
$\begin{split}C(&f_{0},f_{1}|\alpha,\beta)\\ &=\sum_{\epsilon}\left\{\left[f_{0}(\epsilon)-\tilde{P}(0|\epsilon)\right]^{2}+\left[f_{1}(\epsilon)-\tilde{P}(1|\epsilon)\right]^{2}\right\}.\end{split}$
The perturbed probabilities, parameterized by $\alpha$ and $\beta$, were compared to the measured relative frequencies, and the sum of squared residuals across all separations was computed. We then optimized the parameters using nonlinear least-squares minimization. The resulting values characterize the performance of our experimental system and were subsequently used to compute the MLE estimates.
Acknowledgments
We acknowledge J. Szuniewicz for help in developing the EOM control code, R. B. Patel for help in setting up the detectors. We thank B. Brecht, M. G. Raymer, A.M. Steinberg and S. Yu for helpful discussions. This work was supported by the European Union’s Horizon 2020 Research and Innovation Programme Grant No. 899587 Stormytune and the Engineering and Physical Sciences Research Council via the Quantum Computing and Simulation Hub (Grant No. T001062). A. Z. acknowledges a UK Research and Innovation Guarantee Postdoctoral Fellowship under the UK government’s Horizon Europe funding Guarantee (EP/Y029127/1). S.E.T. acknowledges an Imperial College Research Fellowship.
Supplementary Information: Super-resolving frequency measurement with mode-selective quantum memory
IV.1 Fundamental bias at small separations
In the maximum likelihood estimation process, the non-negativity constraint of the parameter space $\Theta$ introduces a positive bias at very small frequency separations, where the distribution of an ideally unbiased estimator may extend into the negative region. Here, we present both an analytical analysis and numerical simulations to characterize this fundamental bias of a two-mode demultiplexing system with crosstalk matrix $\bm{M}$ (Eq. 1 in the main text).
In the absence of crosstalk, the ideal detection probability for mode $n$ is
$P(n|\epsilon)=\frac{\epsilon^{2n}}{16^{n}n!}\exp\left(-\frac{\epsilon^{2}}{16}\right).$ (S1)
And the Fisher information for this HG measurement basis can be calculated as
$\mathcal{F}_{\text{HG}}(\epsilon)=\sum_{n=0}^{\infty}\frac{1}{P(n|\epsilon)}\left(\frac{\partial P(n|\epsilon)}{\partial\epsilon}\right)^{2}.$ (S2)
Considering only the first two modes, the normalized probability of projecting onto the HG1 mode after perturbation by the crosstalk matrix $\bm{M}$ is
$\tilde{P}(1|\epsilon)=\beta-\frac{16(\alpha+\beta-1)}{16+\epsilon^{2}},$ (S3)
and the HG0 projection probability is $\tilde{P}(0|\epsilon)=1-\tilde{P}(1|\epsilon)$. The exact expression for the Fisher information of the non-ideal two-mode projection method can be derived from Eq. S2 in the main text as follows:
$\mathcal{F}(\epsilon)=\frac{(\alpha+\beta-1)^{2}(\epsilon/4)^{2}}{4(1+(\epsilon/4)^{2})^{2}(\alpha+(1-\beta)(\epsilon/4)^{2})(1-\alpha+\beta(\epsilon/4)^{2})}.$ (S4)
We assume the mode crosstalk between the two modes is low, i.e. $\alpha\approx 1$ and $\beta\approx 1$, and the separation is small. Then Eq. S4 can be approximated as Eq. 2 in the main text. Using maximum likelihood estimation (Eq. 3 in the main text), the likelihood function is given by
$\mathcal{L}=N_{0}\ln{\tilde{P}(0|\epsilon)}+N_{1}\ln{\tilde{P}(1|\epsilon)},$ (S5)
where $N_{0}$ and $N_{1}$ represent the experimentally measured detection counts for HG0 and HG1 projections, respectively, with the total count given by $N=N_{0}+N_{1}$. The analytical solution of the MLE estimator can be obtained by solving the critical point equation $d\mathcal{L}/d\epsilon=0$, yielding $\hat{\epsilon}=\sqrt{\frac{16(N_{1}-(1-\alpha)N)}{\beta N-N_{1}}}$. However, this solution is valid only when both the numerator and denominator are non-negative; thus, we require $N_{1}\geq(1-\alpha)N$ (We neglect the condition $N_{1}<\beta N$, as the probability of $N_{1}>\beta N$ is negligible for small separations and low mode crosstalk). This condition corresponds to the range of the modal prediction discussed in the main paper. If a combination of $N_{0}$ and $N_{1}$ values falls outside this range, the likelihood function $\mathcal{L}$ achieves its maximum at the non-differentiable boundary point $\epsilon=0$. In such cases, the MLE defaults to $\hat{\epsilon}=0$. Therefore, the MLE estimators are
$\hat{\epsilon}_{\text{MLE}}=\begin{cases}\sqrt{\frac{16(N_{1}-(1-\alpha)N)}{\beta N-N_{1}}},&\text{if }N_{1}\geq(1-\alpha)N,\\ 0,&\text{if }N_{1}<(1-\alpha)N.\end{cases}$ (S6)
The maximum likelihood estimation method inherently produces only non-negative estimators based on the detected counts. When $N_{0}$ and $N_{1}$ result in a negative estimate, the estimator is set to zero. Consequently, this introduces a bias toward positive values, as all negative estimates are truncated to zero.
Since we employ photon count detection, the detection counts $N_{0}$ and $N_{1}$ follow a Poisson distribution. Specifically, we model $N_{1}$ as $\text{Poisson}(\mu_{1})$, where $\mu_{1}=\tilde{P}(1|\epsilon)N$. Substituting this distribution into the estimator formula in Eq. S6, we derive the mean square (MSE) error as
$\text{MSE}(\epsilon,N)=\sum_{k=0}^{(1-\alpha)N-1}\epsilon^{2}\cdot\frac{\mu_{1}^{k}e^{-\mu_{1}}}{k!}+\sum_{k=(1-\alpha)N}^{\beta N-1}\left(4\sqrt{\frac{k-(1-\alpha)N}{\beta N-k}}-\epsilon\right)^{2}\cdot\frac{\mu_{1}^{k}e^{-\mu_{1}}}{k!}.$ (S7)
(Here, we neglected the $N_{1}>\beta N$ part, as it is extremely unlikely to occur for small separations when $\beta\approx 1$.)
Figure S1: Simulation results of the mean square error (MSE) and the unbiased Cramér-Rao lower bound (CRLB) for maximum likelihood estimation. (A) Simulated MSE, where the white region indicates the onset of truncation due to the non-negativity constraint. (B) MSE with the non-negativity constraint applied. (C) The unbiased CRLB. (D) The difference between the CRLB and the MSE (CRLB - MSE).
The MSE simulated using our experimentally measured crosstalk is shown in Fig. S1 as a function of the separation $\epsilon$ and the number of photons $N$. In panel (A), the white region indicates where the MLE estimators begin to truncate negative estimates, and panel (B) displays the corresponding MSE with this truncation applied. Panel (C) presents the unbiased Cramér-Rao lower bound (CRLB). By taking the difference between the unbiased CRLB and the MSE (CRLB - MSE), we obtain panel (D), which highlights a transition: for large separations, the MSE exceeds the unbiased CRLB (blue regions), but as the separation decreases and with larger bias, it drops below the CRLB (yellow regions).
IV.2 Quantum memory interaction
To characterize the mappings between the signal mode and the spin-wave mode, we can describe the mapping using Green’s functions [55], as the Raman memory equations are linear:
$B_{\mathrm{stor}}(z)=\int_{-\infty}^{+\infty}\mathrm{d}\tau\,K_{1}(z,\tau)S_{\mathrm{in}}(\tau),$ (S8)
$S_{\mathrm{out}}(\tau)=\int_{0}^{L}\mathrm{d}z\,K_{2}(z,\tau)B_{\mathrm{stor}}(z),$ (S9)
where $S_{\text{in}}(\tau)$ is the input signal mode, $B_{\text{stor}}(z)$ is the stored spin-wave mode, $S_{\text{out}}(\tau)$ is the retrieved optical mode. $K_{1}$ and $K_{2}$ are the corresponding Green’s function components that characterize the storage and retrieval processes and contain information about the corresponding control fields, such as temporal modes and intensities.
To derive the memory efficiencies, we first need to define the number of excitations $N$ in each process.
$N_{\text{in(out)}}=\int_{-\infty}^{\infty}d\tau\left<S_{\text{in(out)}}^{\dagger}(\tau)S_{\text{in(out)}}(\tau)\right$> (S10)
$N_{\text{stor}}=\int_{0}^{L}dz\left<B_{\text{stor}}^{\dagger}(z)B_{\text{stor}}(z)\right$>. (S11)
where $N_{\text{in(out)}}$ is the average number of input (output) excitations in the optical field, and $N_{\text{stor}}$ is the average number of the stored spin-wave excitations in the medium. Thus, the storage, retrieval and total memory efficiencies are given as
$\eta_{\mathrm{storage}}$ $=N_{\text{stor}}/N_{\text{in}}=1-N_{\text{tran}}/N_{\text{in}}$ (S12)
$\eta_{\mathrm{retrieval}}$ $=N_{\text{out}}/N_{\text{stor}}=N_{\text{out}}/(N_{\text{in}}-N_{\text{tran}})$
$\eta_{\mathrm{total}}$ $=N_{\text{out}}/N_{\text{in}}.$
where $N_{\text{tran}}=N_{\text{in}}-N_{\text{stor}}$ is the transmitted number of excitations. Experimentally, we used the transmitted photon counts to calculate the storage efficiency, as the number of excitations is not accessible.
The modal properties of the Raman memory can be understood by performing singular value decomposition on the Green functions $K_{1}$ and $K_{2}$ [55]. For the storage process, $K_{1}(z,\tau)$ can be decomposed into
$K_{1}(z,\tau)=\sum_{k}\lambda_{k}\psi_{k}(z)\phi_{k}^{*}(\tau)$ (S13)
where $\lambda_{k}$ are the eigenvalues of the decomposition, sets of $\psi_{k}(z)$ and $\phi_{k}(\tau)$} represent the orthonormal bases of spin-waves and input temporal modes, respectively. In a single-mode quantum memory, only a specific input temporal mode $\phi_{1}(\tau)$ is efficiently coupled and stored as a corresponding spin-wave mode $\psi_{1}(z)$, while all orthogonal modes $\phi_{k}(\tau)$ for $k\neq 1$ pass through the medium unaffected. This selective interaction effectively allows the memory to function as a mode filter, storing one targeted signal while leaving others undisturbed. Crucially, the mode that is stored can be tuned by shaping the control pulse used in the Raman interaction. Adjusting the pulse’s shape, timing, and intensity modifies the Green function $K_{1}(z,\tau)$, which in turn determines the set of temporal modes $\phi_{k}(\tau)$} and their associated coupling strengths, represented by eigenvalues $\lambda_{k}$.
IV.3 Estimation results with different detected photon budgets
Figure S2: MLE estimation results for different total detection counts. (A), (B), and (C) correspond to total detection counts of $N=$ $2\times 10^{3}$, $10\times 10^{3}$ and $100\times 10^{3}$, respectively. The upper panels show the MLE estimators, along with the corresponding quantum bounds (blue region) and direct intensity measurement bounds (gray region). Error bars represent the estimator variances. The lower panels display the MSE (red line), with contributions from bias squared (blue region) and variance (green region).
To compare the performance of MLE estimation under different photon budgets, we present results for total detection counts of $N=$ $2\times 10^{3}$, $10\times 10^{3}$ and $100\times 10^{3}$ in Fig. S2. In all three cases, the MLE estimators closely follow the ground truth (dashed black line). The error bars fall below the intensity detection bound (gray region) but remain above the quantum limit (blue region), demonstrating improved precision over direct intensity (DI) measurements. As expected, the estimator variance decreases with increasing photon counts. The lower panels show the MSE (red line) along with its constituent components, as defined in Eq. 4 in the main text: variance (green area) and bias squared (blue area). The sum of these two contributions aligns well with the overall MSE of the estimators.
IV.4 Different retrieval modes and retrieval times
Figure S3: Performance of MLE estimation at different retrieval times. (A), (B), and (C) correspond to retrieval times of 150 ns, 200 ns, and 250 ns, respectively. In each case, the upper panels show the MLE estimates (red dots), along with the quantum bounds (blue region) and the direct intensity measurement bounds (gray region). The lower panels display the detected signal pulse sequences for two spectral lines with zero separation and zero relative phase. The input signal is shown in gray; the first red pulse indicates leaked signal, and the second red pulse corresponds to the retrieved signal. The read-in control pulse is in the HG0 mode, and the read-out control pulse is in the HG1 mode.
Photonic quantum memories are designed to store photons over user-defined delay times. Raman quantum memories, in particular, offer additional functionality such as temporal mode conversion, where the stored signal can be retrieved into a temporal mode defined by the read-out control pulse. These combined capabilities of storage and mode conversion make Raman memories more versatile than previously demonstrated mode-filtering superresolution schemes, and could enable applications like distributed sensing networks, as discussed in the main text. Here, we demonstrate the superresolution performance of a Raman memory using different retrieval times and output temporal modes.
The maximum storage time is generally limited by thermal diffusion in the warm vapor cell: atoms carrying the spin-wave coherence drift out of the interaction region (defined by the control beam size), rendering the stored signal unrecoverable. The typical coherence time is on the order of microseconds [72]. In our experiments, we characterize the superresolution performance of the Raman memory at storage times of 150 ns, 200 ns, and 250 ns, as shown in Fig. S3. These durations are chosen based on constraints from our delay fiber length and pulse sequence design. For all three retrieval times, the superresolution performance remains comparable, as illustrated in the upper panels of the figure. The lower panels show the raw detection counts from a representative experimental run, where the input signal had zero frequency separation and zero relative phase. The retrieved signals appear as the second red pulse, occurring precisely at 150 ns, 200 ns, and 250 ns, corresponding to retrieval using an HG1 mode as the control read-out pulse.
Figure S4: Performance of MLE estimation with different retrieval modes. (A) and (B) correspond to retrieval control modes of HG0 and HG1, respectively. The upper panels show the MLE estimates (red dots), along with the quantum bounds (blue region) and the direct intensity measurement bounds (gray region). The lower panels display the detected signal pulse sequences for two spectral lines with zero separation and zero relative phase. The input signal is shown in gray; the first red pulse indicates leaked signal, and the second red pulse corresponds to the retrieved signal. The read-in control pulse is in the HG0 mode.
We also performed the experiment using different retrieval modes to demonstrate the flexibility in choosing the read-out control pulse. The results are shown in fig. S4, where the control read-out pulses are set to HG0 and HG1, respectively. The superresolution performance remains comparable in both cases, and the retrieved signals (the second red pulse) clearly exhibit the corresponding temporal mode shapes.
IV.5 Sources of mode crosstalk
In this work, we employed an imperfect two-mode projection method to generate our estimators, incorporating all imperfections into the parameters $\alpha$ and $\beta$ in the perturbation matrix $\bm{M}$ (Eq. 1 in the main text). Below, we discuss the primary sources of crosstalk that contribute to the mode crosstalk and the bias in the raw estimator.
The primary source of temporal mode crosstalk arises from the memory’s storage interaction. Although the memory operates approximately in a single-mode regime, increasing the control field strength induces stronger AC Stark shifts. This leads to the coupling of orthogonal signal modes into the spin wave, distorting the ideal projection probabilities and introducing bias. Fig. S5 illustrates how different storage efficiencies (i.e. different control read-in pulse power) affect the superresolution performance and estimation precision relative to DI methods.
Panel (A) shows the raw estimators $\hat{\epsilon}=4\sqrt{N_{1}/N_{0}}$ for different storage efficiencies. As storage efficiency increases, so does crosstalk, resulting in greater bias in the raw estimators. In panel (B), we apply maximum likelihood estimation and plot the corresponding unbiased CRLB for the two-mode demultiplexing method, alongside the quantum CRLB and the CRLB for DI methods. Higher storage efficiency leads to an increase in the CRLB, resulting in reduced estimation precision. This trend is further quantified in panel (C), where the precision improvement factor, $\mathcal{F}/\mathcal{F}_{\text{DI}}$, is plotted as a function of separation $\epsilon$. As the storage efficiency decreases, the superresolution parameter increases nearly logarithmically. However, very high precision improvement using very low storage efficiencies is impractical in experiments, as fewer signal photons are retrieved, constant background and control field leakage reduce the signal-to-noise ratio, potentially increasing the observed crosstalk and estimation bias.
Another major limiting factor contributing to excess crosstalk in our experiment is control field leakage, which introduces a constant background noise. Due to the small frequency separation (9.2 GHz) between the signal and control fields, completely filtering out the control field is technically challenging. As a result, residual control photons are detected and contribute equally to both HG0 and HG1 projection measurement counts, adding uniform noise across all separations that distorts the projection probabilities.
To quantify this effect, we simulate how constant background counts influence the bias of the raw estimators and the precision improvement of MLE estimators. In this simulation, we fix the storage efficiency at 40% and introduce a constant leakage level ranging from 0 to 2% of the total detected counts. The results are shown in fig. S6. Panel (A) shows that increasing control leakage leads to larger bias in the raw estimator, similar to the effect of increasing storage efficiency. The rate of bias increase slows as the leakage grows uniformly. Panel (B) presents the corresponding precision improvement, which diminishes with higher background noise due to the increasing bias.
Figure S5: Effect of storage efficiency on superresolution performance. (A) Raw estimators $\hat{\epsilon}=4\sqrt{N_{1}/N_{0}}$ for various storage efficiencies, showing increased bias at higher efficiencies due to stronger temporal mode crosstalk. (B) Unbiased CRLB for the two-mode demultiplexing method at different storage efficiencies, compared to the quantum CRLB and the CRLB for the direct intensity measurement method. (C) Precision enhancement $\mathcal{F}/\mathcal{F}_{\text{DI}}$ as a function of separation $\epsilon$, demonstrating logarithmic improvement for small separations.
Figure S6: Effect of control leak counts on estimation bias and precision. (A) The bias in the raw estimators increases with control leakage, expressed as a percentage of the total detection counts. (B) Precision enhancement of the mode filtering method decreases as background noise increases. Storage efficiency is fixed at 40%.
Additional sources of mode crosstalk include detector dark counts, signal background from incomplete extinction of the electro-optic modulator, and four-wave mixing (FWM) noise generated during the Raman interaction. In our experiment, we minimized the signal background using a Pockels cell and characterized both the dark counts and residual background, which were subtracted from the measured data. Furthermore, we applied appropriate detunings to the Raman transition to naturally suppress FWM noise, as described in Ref. [71]. Thus, the contribution of these sources is minimal.
IV.6 Precise pulse carving
Figure S7: Intensity profiles of experimental HG control pulses. (A) and (B) show the measured intensity profiles of the HG0 and HG1 control pulses, respectively. In the upper panels, the experimentally measured normalized profiles (solid red) are compared with the theoretical normalized HG profiles (dashed black). The lower panels display the discrepancies between the experimental and theoretical profiles.
To improve our pulse, carving system, we applied frequency response correction to our electronic system. In a linear time-invariant (LTI) system, where our electronic system can be approximated as such, the electric signal output $h(t)$ of the system to an arbitrary electric signal input function $x(t)$ can be approximately described by a linear response function $R(t-t^{\prime})$ as
$h(t)\approx\int_{-\infty}^{t}dt^{\prime}R(t-t^{\prime})x(t^{\prime}).$ (S14)
In the frequency domain, the frequency response function $\tilde{R}(\omega)$ can be calculated as,
$\tilde{R}(\omega)=\frac{\tilde{h}(\omega)}{\tilde{x}(\omega)}.$ (S15)
Once the response function is known, a target signal $h^{\prime}(t)$ can be precisely generated by using the response-corrected input signal $x^{\prime}(t)$:
$x^{\prime}(t)=\mathcal{F}^{-1}\left[\frac{\mathcal{F}(h^{\prime}(t))}{\tilde{R}(\omega)}\right].$ (S16)
where $\mathcal{F}$ denotes the Fourier transform and $\mathcal{F}^{-1}$ denotes the inverse Fourier transform.
To demonstrate our high-fidelity optical pulse carving system, we present the averaged and normalized intensity profiles of the HG0 and HG1 optical control pulses, as measured by a photodiode and an oscilloscope (1024 averages). fig. S7(A) and (B) show the experimentally measured pulse shapes (red) compared against the theoretical intensity distributions (black dashed lines). The discrepancies between the theoretical and experimental data are plotted in the lower panels. The deviation between the measured and theoretical profiles is small, typically not exceeding 2% relative to the maximum.